### v1

# Universal Predictors for Mixing Time more than Liouvillian Gap

Abstract
We analyze the mixing time of open quantum systems governed by the Lindblad master equation, showing it is not only determined by the Liouvillian gap, but also the trace norm of the lowest excited state of Liouvillian superoperator. By utilizing these universal predictors of mixing time, we establish general conditions for the fast and rapid mixing respectively. Specifically, we derive rapid mixing conditions for both the strong and weak dissipation regimes, formulated as sparsity constraints on the Hamiltonian and the local Lindblad operators. Our findings provide a general framework for calculating mixing time and offer a guide for designing dissipation to achieve desired mixing speeds, which has significant implications for efficient experimental state preparation.
Introduction— The study of open quantum systems is of paramount importance, both theoretically and experimentally [10, 24]. In real-world experiments, the interaction between a quantum system and its environment is almost inevitable. Therefore, it is crucial to model quantum many-body systems as open systems and understand their dynamics, which are closely related to phenomena like information scrambling [2, 66, 56, 39, 70, 7, 63, 50], quantum chaos [65, 53, 41, 67, 30, 78, 75], and quantum phase transitions [43, 12, 61, 62, 21, 77]. Furthermore, the dynamics of open systems can be harnessed for dissipative state preparation [60, 59, 14, 37, 71, 38, 16, 6, 34, 5, 48, 64, 44, 17, 74, 47, 51], where carefully designed dissipation acts as a tool to steer a system toward a specific initial state relevant for quantum computation and quantum simulation [22, 52, 25, 8, 27, 58, 40, 55].
A key concept in the study of open system dynamics is the mixing time, defined as the minimum time it takes for an arbitrary initial state to evolve to the system’s steady state [54, 79]. In dissipative state preparation, mixing time is a vital metric because it quantifies the practical time required to prepare a desired state, thus determining whether a protocol is experimentally feasible [18, 74]. Also, an estimation of the mixing time typically plays a crucial role in many of the quantum algorithmic applications to yield efficient quantum algorithms [28, 1]. Also, mixing time is related to entropy [31, 46, 13], entanglement dynamics, and the structure of the Lindblad spectrum and the dissipative phase transition, where at the transition point, the mixing time diverges [46, 79, 45, 42, 57, 33, 32].
Lots of progress has been made in analyzing the mixing time for the open system dynamics. Previous studies have shown that for Lindbladians satisfying quantum detailed balance, the system’s dynamics can be mapped to a local, frustration-free parent Hamiltonian in a doubled Hilbert space. The spectral gap of this Hamiltonian then provides a useful lower bound for the Lindbladian’s spectral gap, offering a powerful tool for estimating mixing time [15, 68, 35]. A different approach uses the modified logarithmic Sobolev inequality [73, 9], which analyzes the exponential decay of relative entropy. This method can provide a tighter, polylog($N$) mixing time bound, offering a significant improvement over bounds based solely on the spectral gap [36, 11, 4, 3, 23]. Besides, other research has estimated mixing time by analyzing the Hamiltonian and dissipative parts separately, although this is applicable in limited cases [20]. Despite the recent progress, obtaining the mixing time of the Lindbladian dynamics, and analyzing its dependent of system size is typically difficult and can sometimes be the bottleneck of the analysis.
In this letter, we map the definition of mixing time of the open quantum system to the distance between wavefunctions in a doubled Hilbert space, directly connecting its definition to the Liouvillian gap and the trace norm of the lowest excited state. Using these two quantities as universal predictors, we identify general conditions for fast mixing and rapid mixing. We further apply our framework to the weak- and strong-dissipation regimes, where the resulting rapid mixing criteria can be expressed as sparsity constraints on the Hamiltonian and the local Lindblad operators. Overall, our results provide a general route to estimating mixing time and offer practical principles for engineering open-system dynamics for efficient state preparation.
Mixing Time and the Lindblad Spectrum—We consider the open system whose time evolution is driven by the Lindblad Master equation [26, 49]
$\frac{\partial{\rho}}{\partial t}=-i[H,\rho]+2\gamma\sum_{\alpha}K_{\alpha}\rho K^{\dagger}_{\alpha}-\gamma\sum_{\alpha}\{K^{\dagger}_{\alpha}K_{\alpha},\rho\}.$ (1)
Here, $\gamma$ is the dissipation strength, and $K_{\alpha}$ is the Lindblad operator. Below, we use $e^{\mathcal{L}t}[\rho_{0}]$ to denote the density matrix at time $t$, given an initial density matrix $\rho_{0}$ driven by the Lindblad master equation.
We recall the definition of the mixing time, which quantifies how long it takes for an arbitrary initial state to evolve close to the steady state $\sigma$ of the dynamics. Closeness is measured by the trace distance
$D(\rho,\sigma)=\frac{1}{2}\|\rho-\sigma\|_{1},$ (2)
where the trace norm is $\|A\|_{1}\equiv\mathrm{Tr}\!\left[\sqrt{A^{\dagger}A}\right]$. For Hermitian $A$, $\|A\|_{1}=\mathrm{Tr}|A|$.
The mixing time at accuracy $\eta$ is then defined as
$\tau_{\mathrm{mix}}(\eta)=\min\left\{t\,\big|\,D\!\left(e^{\mathcal{L}t}[\rho_{0}],\sigma\right)\leq\eta,\ \forall\,\rho_{0}\right\}.$ (3)
That is, $\tau_{\mathrm{mix}}(\eta)$ is the minimal time such that, for arbitrary initial state $\rho_{0}$, the evolved state $e^{\mathcal{L}t}[\rho_{0}]$ is within trace distance $\eta$ of the steady state $\sigma$. Since density matrices are Hermitian, this distance can also be written as
$D(\rho(t),\sigma)=\frac{1}{2}\,\mathrm{Tr}\,|\rho(t)-\sigma|.$ (4)
We now introduce a mapping that relates the mixing time directly to spectral properties of the Lindbladian. Using vectorization (the Choi–Jamiołkowski isomorphism) [69, 80], we map the density matrix to a state in a doubled Hilbert space,
$\begin{split}&\rho=\sum_{m,n}\rho_{mn}\,|m\rangle\langle n|\\ \ \longmapsto\ &|\psi_{\rho}^{D}\rangle=\sum_{m,n}\rho_{mn}\,|m\rangle_{L}\otimes|n\rangle_{R}.\end{split}$ (5)
Here $L$ and $R$ denote the left and right copies, respectively. Under this mapping, the Lindblad master equation becomes a Schrödinger-like equation in doubled space [76],
$i\,\partial_{t}|\psi_{\rho}^{D}(t)\rangle=H^{D}\,|\psi_{\rho}^{D}(t)\rangle,$ (6)
with a generally non-Hermitian generator $H^{D}=H_{s}-iH_{d}$ given by
$\begin{split}H_{s}=&H_{L}\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes H_{R}^{T},\\ H_{d}=&\gamma\sum_{\alpha}\Big[-2\,K_{\alpha,L}\otimes K_{\alpha,R}^{*}\\ &+(K_{\alpha}^{\dagger}K_{\alpha})_{L}\otimes\mathcal{I}_{R}+\mathcal{I}_{L}\otimes(K_{\alpha}^{\dagger}K_{\alpha})_{R}^{*}\Big].\end{split}$ (7)
Here $H_{L/R}$ acts as $H$ on the corresponding copy, $T$ denotes transpose, denotes complex conjugation, and $\mathcal{I}$ denotes the identity.
Expanding the initial state in right eigenstates of $H^{D}$,
$|\psi_{\rho_{0}}^{D}\rangle=\sum_{j}c_{j}\,|\psi_{\sigma_{j}}\rangle,$ (8)
the time evolution is
$|\psi_{\rho_{0}}^{D}(t)\rangle=\sum_{j}c_{j}\,e^{-i\epsilon_{j}t}\,|\psi_{\sigma_{j}}\rangle=\sum_{j}c_{j}\,e^{-i\alpha_{j}t-\beta_{j}t}\,|\psi_{\sigma_{j}}\rangle,$ (9)
where $\epsilon_{j}=\alpha_{j}-i\beta_{j}$ are the eigenvalues of $H^{D}$ (the Lindblad spectrum). The real parts $\alpha_{j}$ generate oscillations, while $\beta_{j}\geq 0$ set decay rates. [^1] Complete positivity implies $\beta_{j}\geq 0$. Trace preservation fixes the steady mode: $c_{0}=1$, $\mathrm{Tr}\,\sigma_{0}=1$, and $\mathrm{Tr}\,\sigma_{j}=0$ for $j\neq 0$. Hermiticity preservation implies $\sigma_{j}=\sigma_{j}^{\dagger}$.
Consequently, the deviation from the steady state $\sigma_{0}$ is
$|\psi_{\rho}^{D}(t)\rangle-|\psi_{\sigma_{0}}\rangle=\sum_{j\neq 0}c_{j}\,e^{-i\alpha_{j}t-\beta_{j}t}\,|\psi_{\sigma_{j}}\rangle,$ (10)
where $|\psi_{\sigma_{0}}\rangle$ is the steady-state mode and $\{|\psi_{\sigma_{j}}\rangle\}_{j>0}$ are the decaying modes.
The mixing time, as defined in Eq. (3), is the minimum time required for an arbitrary initial state to converge to the steady state. This mixing time is ultimately determined by the slowest decaying mode, which corresponds to the Liouvillian’s smallest non-zero eigenvalue in terms of its imaginary part. The magnitude of this eigenvalue is known as the Liouvillian gap $\Delta$. Therefore, the worst case for the mixing time is obtained when the initial state has a non-zero overlap with the eigenmode associated with this gap. We denote this slowest decaying eigenmode as $|\psi_{\sigma_{1}}\rangle$.
A convenient initial state with nonzero overlap on the slowest mode is
$|\psi_{\rho_{0}}\rangle=|\psi_{\sigma_{0}}\rangle+c_{1}|{\psi_{\sigma_{1}}}\rangle,$ (11)
where $\sigma_{0}$ is the steady state and $\sigma_{1}$ is the slowest decaying mode of the Lindbladian with eigenvalue $\Delta+i\alpha_{1}$ ($\Delta>0$). One might wonder whether the coefficient $c_{1}$ could grow with system size. This is ruled out by a purity bound: under vectorization, $\|\lvert\psi_{\rho_{0}}\rangle\|^{2}=\mathrm{Tr}(\rho_{0}^{2})\leq 1$, so the overlap (and hence $c_{1}$) is at most $O(1)$. We may therefore treat $c_{1}$ as an $O(1)$ constant.
At time $t$,
$|\psi_{\rho_{0}}(t)\rangle-|\psi_{\sigma_{0}}\rangle=c_{1}e^{-(\Delta-i\alpha_{1})t}\,|\psi_{\sigma_{1}}\rangle.$ (12)
Consequently, the trace distance to the steady state is
$D(\rho(t),\sigma_{0})=\frac{|c_{1}|}{2}\operatorname{Tr}|e^{-i\alpha_{1}t-\Delta t}\sigma_{1}|=\frac{|c_{1}|}{2}e^{-\Delta t}\operatorname{Tr}|\sigma_{1}|.$ (13)
The mixing time is thus
$\tau_{\textit{mix}}(\eta)=\Delta^{-1}\left[\log(\operatorname{Tr}|\sigma_{1}|)-\log{(\frac{2\eta}{|c_{1}|})}\right].$ (14)
This reformulation of the mixing time indicates that the mixing time of an open system is set by the inverse Liouvillian gap, multiplied by a logarithmic factor controlled by the trace norm of the lowest excited state of the doubled-space Hamiltonian. One might worry that higher excited states, despite their larger decay rates, could have sufficiently smaller trace norms to yield a larger overall mixing-time bound. We explain in Appendix C why it is nevertheless sufficient to focus on the lowest excited state  [^2] We focus on the lowest excited state for two reasons. First, the fast mixing condition imposes no explicit constraints on trace norms across the spectrum, so the lowest excited state provides a natural and minimal diagnostic. Second, in both the weak- and strong-dissipation regimes, the rapid mixing requirement on the trace norm takes the same functional form for all excited states, and is therefore not specific to the lowest excited state. See Appendix C for details..
The trace norm of the lowest excited state has some explicit constraints. Using the Cauchy-Schwarz inequality, we have:
$\begin{split}\operatorname{Tr}|\sigma_{1}|&=\sum_{j=1}^{N}|\lambda_{j}|\leq\sqrt{(\sum_{j=1}^{N}\lambda_{j}^{2})(\sum_{j=1}^{N}1)}=\sqrt{N}.\end{split}$ (15)
Here, $N$ is the dimension of the open system’s Hilbert space. The final equality holds because $\operatorname{Tr}(\sigma_{1}^{2})=1$, which is due to the normalization of the eigenstates. Thus,
$D(e^{\mathcal{L}t}\rho_{0},\sigma)\leq\frac{|c_{1}|}{2}\sqrt{N}e^{-\Delta t}.$ (16)
The equality is taken when $\sigma_{1}$ has eigenvalues that are pairs of $(\frac{1}{\sqrt{N}},-\frac{1}{\sqrt{N}})$ [^3] Recall that the density matrix remains trace equal to 1 along the time evolution, thus $\operatorname{Tr}(\sigma_{j})=0,\forall j\neq 0$.. Combining this with the definition of mixing time, we arrive at a general upper bound of mixing time:
$\tau_{\textit{mix}}(\eta)\leq\frac{1}{2}\Delta^{-1}\left[\log(N)-2\log{(\frac{2\eta}{|c_{1}|})}\right].$ (17)
For a system of size $L$ and local Hilbet space dimension $d_{0}$,
$\tau_{\text{mix}}(\eta)\leq\tfrac{1}{2}\Delta^{-1}\left[L\log d_{0}-2\log{(\frac{2\eta}{|c_{1}|})}\right].$ (18)
This bound of mixing time is general but not always saturated. For example, non-Hermitian skin effects can localize eigenstates near boundaries [29], reducing effective support and giving $\tau_{\text{mix}}\sim\Delta^{-1}\text{poly}(\log L)$ for total system size $L$. Mixing time is therefore governed jointly by the Liouvillian gap and the spatial structure of eigenstates.
Requirements for fast and rapid mixing.— We first recall two standard notions of efficient convergence to the steady state. The dynamics is fast mixing if the mixing time grows at most polynomially with the system size, $\tau_{\mathrm{mix}}=\mathcal{O}(\mathrm{poly}(L))$, and rapid mixing if it grows at most polylogarithmically, $\tau_{\mathrm{mix}}=\mathcal{O}(\mathrm{poly}[\log L])$. Our general bound in Eq.(14) immediately translates these notions into requirements on the inverse Liouvillian gap $\Delta^{-1}$ and on the trace norm of the lowest excited state $\mathrm{Tr}|\sigma_{1}|$.
Theorem 1 (Fast-mixing condition).
Fast mixing follows if the inverse Liouvillian gap is polynomially bounded in the system size $L$,
$\Delta^{-1}\leq\mathcal{O}\!\bigl(\mathrm{poly}(L)\bigr).$ (19)
The trace norm of the lowest excited state is not restrictive for fast mixing: since $\sigma_{1}$ is an operator on a Hilbert space of dimension $d_{0}^{L}$, from Eq.(15), it is already polynomial in $L$. Representative examples where $\Delta^{-1}$ is polynomial include dissipative preparation of stabilizer states [72] and Davies generators for high-temperature commuting local Hamiltonians [19].
Rapid mixing imposes stronger constraints. In addition to a polylogarithmic gap scaling, one must also prevent $\mathrm{Tr}|\sigma_{1}|$ from becoming exponentially large in $L$.
Theorem 2 (Rapid-mixing conditions).
Rapid mixing holds if both of the following conditions are satisfied:
  1. 1.
    the inverse gap scales at most polylogarithmically,
    $\Delta^{-1}\leq\mathcal{O}\!\bigl(\mathrm{poly}[\log(L)]\bigr);$ (20)
  2. 2.
    the lowest excited state has at most polynomial trace norm,
    $\mathrm{Tr}|\sigma_{1}|\leq\mathcal{O}\!\bigl(\mathrm{poly}(L)\bigr).$ (21)
Unlike fast mixing, rapid mixing requires an explicit constraint on $\sigma_{1}$. A simple sufficient mechanism is a global $U(1)$ symmetry shared by both the Hamiltonian and the dissipators: the Liouvillian then decomposes into charge sectors, and if the relevant sectors have Hilbert-space dimension growing at most polynomially with $L$, the corresponding operator norms and trace norms remain polynomially bounded.
For convenience, we summarize the general requirements for fast and rapid mixing in Table 1.
$\tau_{\mathrm{mix}}$ $\Delta^{-1}$ $\mathrm{Tr}|\sigma_{1}|$
fast mixing $\leq\mathrm{poly}(L)$ no constraint
rapid mixing $\leq\mathrm{poly}[\log(L)]$ $\leq\mathrm{poly}(L)$
Table 1: Conditions for fast and rapid mixing in terms of the inverse Liouvillian gap $\Delta^{-1}$ and the trace norm of the lowest excited state $\mathrm{Tr}|\sigma_{1}|$.
We next specialize these general criteria to the weak- and strong-dissipation regimes, which yield concrete model-dependent sufficient conditions. The resulting rapid mixing conditions are summarized in Table 2.
dissipation Hamiltonian $H$ Lindblad operator $K$
weak $H$ gapped $\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|E_{s}\rangle,\ N\!\bigl(\alpha,K,|E_{s}\rangle\bigr)\leq\mathrm{poly}(L)$
strong $\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|\epsilon_{s}\rangle,\ N\!\bigl(\alpha,H,|\epsilon_{s}\rangle\bigr)\leq\mathrm{poly}(L)$ boundary dissipation
Table 2: Sufficient conditions for rapid mixing in the weak- and strong-dissipation regimes, expressed as sparsity constraints on the Hamiltonian $H$ and the local Lindblad operator $K$. For a fixed eigenstate $|\epsilon_{s}\rangle$ of $\{K_{j}\}$, $N(\alpha,H,|\epsilon_{s}\rangle)$ counts the number of indices $p$ with $|H_{sp}|\geq e^{-\alpha L}$. For a fixed eigenstate $|E_{s}\rangle$ of $H$, $N(\alpha,K,|E_{s}\rangle)$ counts the number of indices $p$ with $|K_{sp}|\geq e^{-\alpha L}$.
Strong-dissipation regime— In this regime, the dissipative part of the Liouvillian dominates, and the Hamiltonian (Schrödinger) contribution can be treated perturbatively. We use degenerate perturbation theory to estimate the Liouvillian gap and the mixing time, and obtain the following fast-mixing corollary:
Corollary 3.
In a one-dimensional open quantum system governed by a Lindblad equation, strong boundary dissipation implies fast mixing whenever the Lindblad operators do not commute with the system Hamiltonian.
For boundary dissipation, the unperturbed Liouvillian has a highly degenerate steady-state manifold: because the dissipator acts only at the edges, changes in bulk degrees of freedom do not shift the zeroth-order eigenvalue. A Hamiltonian perturbation $V^{D}$ therefore has two distinct effects: it (i) mixes different zeroth-order steady states via bulk terms, and (ii) couples the steady manifold to decaying modes through boundary terms. We characterize the corresponding matrix elements by typical scales,
$\langle\psi_{0k}|V^{D}|\psi_{1\alpha}\rangle\sim J,\qquad\langle\psi_{0k}|V^{D}|\psi_{0l}\rangle\sim J^{\prime},$
where $\{|\psi_{0k}\rangle\}$ spans the degenerate zeroth-order manifold (chosen to diagonalize $V^{D}$ within that subspace) and $\{|\psi_{1\alpha}\rangle\}$ denotes zeroth-order excited modes.
The leading nonzero decay rate of the perturbed steady manifold arises at second order via virtual transitions into the excited sector. Standard degenerate perturbation theory gives the effective eigenvalue
$E_{0k}\approx E^{(0)}_{0k}+V^{D}_{0k,0k}+\sum_{q}\frac{|V^{D}_{0k,1q}|^{2}}{E^{(0)}_{0k}-E^{(0)}_{1q}}$. (22)
The diagonal shift $V^{D}_{0k,0k}$ contributes only to the real part, while the Liouvillian gap is set by the dissipative contribution from the second-order term,
$\Delta\equiv\mathrm{Im}\,E_{0k}\sim\sum_{q}\frac{|V^{D}_{0k,1q}|^{2}}{E^{(0)}_{1q}-E^{(0)}_{0k}}\sim c\,\frac{J^{2}}{\gamma_{\mathrm{loc}}},$ (23)
with an $\mathcal{O}(1)$ constant $c$ (for our boundary coupling structure, c=8c=8). [^4] The factor $8$ counts: (i) action on the left or right copy, (ii) action at the left or right boundary, and (iii) raising or lowering the boundary state, giving $2^{3}=8$ possibilities. Crucially, $\Delta$ does not shrink with $L$: only Hamiltonian terms involving boundary sites connect the steady manifold to decaying modes, so the number of contributing excited states is set by the boundary size, i.e., $\mathcal{O}(1)$ in 1D.
Using $\tau_{\mathrm{mix}}(\eta)\sim\Delta^{-1}\log\!\bigl(\mathrm{Tr}|\sigma_{1}|\bigr)$ (up to the standard $\eta$-dependent additive constant), we obtain
$\tau_{\mathrm{mix}}(\eta)\sim\frac{\gamma_{\mathrm{loc}}}{c\,J^{2}}\log\!\bigl(\mathrm{Tr}|\sigma_{1}|\bigr).$ (24)
Moreover, the first-order correction to the lowest excited mode generically spreads over the entire degenerate zeroth-order manifold. Since that manifold contains $d_{0}^{2L-2}$ basis states,
$\log\!\bigl(\mathrm{Tr}|\sigma_{1}|\bigr)\leq\log\!\bigl(\sqrt{d_{0}^{2L-2}}\bigr)=(L-1)\log d_{0},$
and this bound is saturated here because the $V^{D}$-diagonal basis typically forms superpositions across the full manifold. Therefore,
$\tau_{\mathrm{mix}}(\eta)\sim\frac{\gamma_{\mathrm{loc}}}{c\,J^{2}}(\log d_{0})\,L,$ (25)
which is polynomial in $L$ and hence implies fast mixing. In the Appendix we present the analogous analysis for strong bulk dissipation and show that this perturbative mechanism does not apply there.
Physically, boundary-driven fast mixing is a quantum-Zeno-type stabilization: strong edge dissipation sets an $\mathcal{O}(1)$ scale, while boundary Hamiltonian couplings generate a size-independent gap $\Delta\propto J^{2}/\gamma_{\mathrm{loc}}$. Thus, boundary dissipation can guarantee scalable mixing even for complex bulk dynamics, whereas bulk dissipation typically requires different control assumptions (see Appendix).
Finally, in the strong-dissipation regime, rapid mixing also requires control of the trace norm $\textrm{Tr}|\sigma_{1}|$, as stated in Theorem 2. To make this requirement explicit, we compute the leading correction to lowest excited doubled-space eigenstate in the vectorized $K$-eigenbasis $\{|\epsilon_{p},\epsilon_{q}\rangle\}$,
$|\epsilon_{m},\epsilon_{m}^{(1)}\rangle=-\sum_{p}\frac{H_{pm}}{\epsilon^{(0)}_{pm}}\,|\epsilon_{p},\epsilon_{m}^{(0)}\rangle+\sum_{q}\frac{H_{qm}}{\epsilon^{(0)}_{mq}}\,|\epsilon_{m},\epsilon_{q}^{(0)}\rangle,$ (26)
where $|\epsilon_{m}\rangle$ is the eigenstate of Lindblad operator $K$, $H_{pm}\equiv\langle\epsilon_{p}|H|\epsilon_{m}\rangle$, and $\epsilon^{(0)}_{pm}$ are the corresponding unperturbed doubled-space energy denominators. Requiring $\textrm{Tr}|\sigma_{1}|\leq\mathrm{poly}(L)$ then translates into a sparsity condition on $H$ in the $\{K_{j}\}$ eigenbasis:
$\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|\epsilon_{s}\rangle,\qquad N(\alpha,H,|\epsilon_{s}\rangle)\leq\mathrm{poly}(L).$ (27)
Here $N(\alpha,H,|\epsilon_{s}\rangle)$ counts the number of indices $p$ for which $|H_{sp}|\geq e^{-\alpha L}$, with $|\epsilon_{s}\rangle$ a fixed eigenstate of $\{K_{j}\}$. The proof is given in the Appendix.
Weak-dissipation regime.— In the weak-dissipation regime we treat the dissipator as a perturbation to the unitary (Schrödinger) dynamics. We estimate (i) the Liouvillian gap from the leading dissipative correction to the unperturbed spectrum and (ii) the trace norm of the lowest excited state from the corresponding first-order correction to the eigenvectors. Combining these ingredients yields the mixing-time scaling.
Liouvillian gap. Consider an unperturbed doubled-space eigenstate
$|E_{m},E_{n}^{(0)}\rangle\equiv|E_{m}\rangle_{L}\otimes|E_{n}\rangle_{R},\qquad\epsilon_{mn}^{(0)}=E_{m}-E_{n},$
where $\{|E_{m}\rangle\}$ are eigenstates of $H$. The leading (first-order) correction from the dissipative term is
$\langle E_{m},E_{n}^{(0)}|H_{d}|E_{m},E_{n}^{(0)}\rangle=-i\gamma\Bigl[\delta K_{m}^{2}+\delta K_{n}^{2}+\bigl(\overline{K_{m}}-\overline{K_{n}}\bigr)^{2}\Bigr],$ (28)
with $\overline{K_{m}}\equiv\langle E_{m}|K|E_{m}\rangle$ and
$\delta K_{m}^{2}\equiv\langle E_{m}|K^{2}|E_{m}\rangle-\overline{K_{m}}^{\,2}.$
When $K,H]\neq 0$, typical fluctuations scale as $\delta K_{m}^{2}\sim L^{a}$ with $a\geq 0$. For gapped systems one expects $a=0$, giving a Liouvillian gap of order $\mathcal{O}(\gamma)$; for gapless systems ($a>0$), enhanced fluctuations can parametrically reduce the gap. [^5] If $[K,H]=0$, then $\delta K_{m}^{2}=0$ and decay rates are set by the smallest nonzero value of $(\overline{K_{m}}-\overline{K_{n}})^{2}$. For local $K$ in an ETH-satisfying system this scale is typically $\mathcal{O}(1)$, yielding a size-independent gap. Therefore, a sufficient condition for rapid mixing based on the Liouvillian gap is that $H$ is gapped, which ensures a constant gap.
Trace norm of the lowest excited state. The first-order correction to the corresponding excited eigenvector takes the form
$\begin{split}|E_{m},E_{n}^{(1)}\rangle=-i\gamma\sum_{p,q,s,t}\frac{\bigl(K_{ps}-K_{qt}\bigr)\bigl(K_{sm}-K_{tn}\bigr)}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle,\end{split}$ (29)
with $K_{ps}\equiv\langle E_{p}|K|E_{s}\rangle$. To ensure that the trace norm of the lowest excited state grows at most polynomially with system size, it is sufficient to impose a sparsity constraint on the matrix elements of $K$ in the $\{|E_{s}\rangle\}$ basis:
$\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|E_{s}\rangle,\qquad N(\alpha,K,|E_{s}\rangle)\leq\mathrm{poly}(L),$ (30)
where $N(\alpha,K,|E_{s}\rangle)$ counts the number of indices $p$ such that
$|K_{sp}|\geq e^{-\alpha L}\quad\text{for fixed }|E_{s}\rangle.$
A simple limiting example is the nearest-neighbor quantum Ising model (with Hamiltonian containing only $\sigma_{j}^{z}\sigma_{j+1}^{z}$ terms), with Lindblad operators taken to be arbitrary single-site operators. In this case the Hamiltonian eigenbasis is a product basis, and each local $K$ connects only $\mathcal{O}(1)$ configurations. Consequently one may take $N(\alpha,K,|E_{s}\rangle)=\mathcal{O}(1)$ (e.g., $N=1$ in the strictly diagonal case), illustrating the required sparsity.
Discussion— By using the double-space mapping, we directly connected the mixing time to the Liouvillian gap and the trace norm of the lowest excited states. This connection provides a clear requirement for achieving fast and rapid mixing. This framework also explains previous findings that mixing time in open systems depends on both the Liouvillian gap and the localization length of the eigenstates [29]. Our analysis shows that this is because localization properties directly affect the trace norm of the eigenstates, which in turn influences the mixing time.
We analyzed the mixing time in both the strong and weak dissipation regimes, where the Liouvillian gap and trace norm can be calculated straightforwardly using perturbation theory. This analysis not only simplifies the problem but also reveals general conditions for fast and rapid mixing. These findings are crucial for assessing the efficiency of quantum algorithms and state preparation using dissipative evolution. The general conditions we identified can be used to design specific forms of the Lindblad operator to more efficiently prepare desired quantum states.
This work opens several interesting avenues for future research. First, it is important to compute the Liouvillian gap and the trace norm of the lowest excited states beyond the weak- and strong-dissipation limits, where our perturbative approach breaks down. Second, in connection with rapid mixing, it would be useful to identify conditions under which the trace norm of the lowest excited states is at most polynomial in system size, and to clarify how such conditions relate to features such as the locality of dissipation or the correlation length. A further question is whether changes in the scaling of the Liouvillian gap are tied to changes in the scaling of the lowest-mode trace norm. By tuning model parameters, one can compare the crossover where the inverse Liouvillian gap changes its system-size dependence (e.g., from polynomial to logarithmic) with the crossover where the trace norm changes (e.g., from exponential to polynomial). Determining whether these crossovers coincide—or how far apart they occur—could provide a new way to distinguish relaxation mechanisms. The boundary-driven transverse-field Ising chain, with dissipation only at the ends, provides a concrete setting in which both scalings vary with dissipation strength and can be directly compared. Finally, for dissipative state preparation, rapid mixing is essential to reach the target state on experimentally relevant timescales; the sparsity constraints derived here offer a practical guide for designing protocols that remain experimentally feasible.
Acknowledgements— We thank Jerome Lloyd, Tian-Gang Zhou and Pengfei Zhang for helpful discussions.
Appendix A Appendix A: Mixing time analysis for strong bulk dissipation
In this Appendix, we present a detailed analysis of the mixing time in the strong bulk-dissipation regime. In particular, we show that the perturbative approach breaks down for bulk dissipation in the thermodynamic limit.
We begin by specifying the zeroth-order eigenstates and eigenvalues of the doubled-space (vectorized) Hamiltonian. For simplicity, we take the Lindblad operator to be Hermitian. The corresponding zeroth-order generator can then be written as
$H^{0,D}=-i\gamma\sum_{\alpha=1}^{L}(K_{\alpha,L}\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes K_{\alpha,R}^{*})^{2}.$ (31)
The ground state(GS) of the zeroth-order Hamiltonian is the state that satisfies
$K_{\alpha,L}|\psi^{0}_{GS}\rangle=K_{\alpha,R}^{*}|\psi^{0}_{GS}\rangle$ (32)
for each Lindblad operator. If we choose $\{|m\rangle\}$ as a set of complete orthogonal eigenbases for each local site, then the GS can be constructed as
$|\psi^{0}_{GS}\rangle_{\mathbf{m}}^{D}=\prod_{x=1}^{L}|m_{x}\rangle_{x,L}\otimes|m_{x}\rangle_{x,R}.$ (33)
Here, $\mathbf{m}=(m_{1},m_{2},...,m_{L})$ with $m_{x}=1,2,...,d_{0}$, and $d_{0}$ is the total local Hilbert space dimension. The GS energy is $E_{0}^{(0)}=0$ and GS degeneracy is $d_{0}^{L}$.
We consider the general case that the Hamiltonian does not commute with the Lindblad operators, which means the perturbation can connect the GS to excited states. Since the Hamiltonian acts on all sites, the number of excited states that can be connected to a given ground state by the perturbation is proportional to the total system size, $L$. We consider the equal level spacing of local sites, and we denote the level spacing as $1$. For simplicity, we consider the general case where the Hamiltonian does not commute with the jump operators and the Hamiltonian only connects the nearest energy level. Then, the first excited state can be represented as the combination of the below four states (this combination has to make sure that this excited state after the inverse double space mapping gives a tracelss hermian matrix)
$|\psi^{0}_{x_{0},e,L\pm}\rangle=|m_{x_{0}}\pm 1\rangle_{x_{0},L}\otimes|m_{x_{0},R}\rangle_{x_{0}}\prod_{x\neq x_{0}}|m_{x}\rangle_{x,L}\otimes|m_{x}\rangle_{x,R},$ (34)
and
$|\psi^{0}_{x_{0},e,R\pm}\rangle=|m_{x_{0}}\rangle_{x_{0},L}\otimes|m_{x_{0}}\pm 1\rangle_{x_{0},R}\prod_{x\neq x_{0}}|m_{x}\rangle_{x,L}\otimes|m_{x}\rangle_{x,R}.$ (35)
Since $x_{0}=1,2,...,L$, there are in total $4L$ degenerated first excited states with eigen-energy $E_{1}^{0}=-i\gamma$.
Using the degenerated perturbation theory, the first-order correction to the GS energy is
$E^{(1)}_{0k}=E^{(0)}_{0k}+V^{D}_{0k,0k}+\sum_{q}\frac{|V^{D}_{0k,1q}|^{2}}{E^{(0)}_{0k}-E^{(0)}_{1q}}$. (36)
Here, $0k$ denotes the label of the zeroth degenerate GS basis set that diagonalizes the $V^{D}$. $1q$ denotes the label of $q$-th zeroth excited state.
We assume that the Hamiltonian does not commute with the Lindblad operators, which means the perturbation can connect the ground state to excited states. Since the Hamiltonian acts on all sites, the number of excited states that can be connected to any given ground state by the perturbation is proportional to the total system size, $L$.
We denote the matrix elements of the perturbation as $\langle\psi_{0k}|V^{D}|\psi_{1q}\rangle=J$, where $|\psi_{0k}\rangle$ represents a GS with label $k$ and $|\psi_{1q}\rangle$ represents a first excited state with label $q$. Since the diagonal (Hermitian) shift $V^{D}_{0k,0k}$ only affects the real part, it does not affect the Liouvillian gap. The only contribution to the Liouvillian gap is given by the last term in the Eq.(36)
$\Delta\propto-iL\frac{J^{2}}{\gamma}.$
Because this energy correction diverges as $L\to\infty$, this perturbation analysis is not applicable in the thermodynamic limit.
Appendix B Appendix B: Reason for choosing the lowest excited state in the rewriting of mixing time
In this appendix, we provide the reason for choosing the initial density matrix as a combination of the steady state and the lowest exited state to get an estimation of the mixing time. As we derived in the main text, we arrive at:
$\tau_{\textit{mix}}(\eta)=\Delta^{-1}\left[\log(\operatorname{Tr}|\sigma_{1}|)-\log{(\frac{2\eta}{|c_{1}|})}\right].$ (37)
Here, $\Delta$ is the Liouvillian gap and $\sigma_{1}$ denotes an lowest excited eigenmode (written as an operator in the original Hilbert space, or equivalently as a vector in the doubled space). This reformulation shows that the mixing time of an open system can be expressed as the inverse Liouvillian gap, multiplied by a logarithmic factor set by the trace norm of an excited state in the doubled (vectorized) space.
A potential concern is that higher excited states, while decaying faster, might have substantially smaller trace norms and could in principle yield a larger estimate of the mixing time. For example, one could imagine that for some mode $(\beta_{2},\sigma_{2})$,
$\beta_{2}^{-1}\log(\mathrm{Tr}|\sigma_{2}|)\;>\;\Delta^{-1}\log(\mathrm{Tr}|\sigma_{1}|),$ (38)
so that including $\sigma_{2}$ would appear to increase the bound. Nevertheless, throughout this work we use the lowest excited state as a representative, for the following reasons.
First, the criterion for fast mixing depends only on the Liouvillian gap and is insensitive to the trace norm prefactor. Therefore, when the goal is simply to determine whether the dynamics is fast mixing, the choice of which excited state enters the logarithmic factor is immaterial. For notational simplicity, we refer to the mode used in the trace-norm factor as the lowest excited state.
Second, in the perturbative regimes where we derive explicit constraints, the structure of the eigenstates is uniform across the spectrum, so that the trace-norm constraints relevant for rapid mixing are not special to the lowest excited sector. To illustrate this in the weak-dissipation limit, the first-order correction to an arbitrary unperturbed eigenstate $|E_{m},E_{n}^{(0)}\rangle$ takes the form
$\begin{split}|E_{m},E_{n}^{(1)}\rangle&=-i\gamma\sum_{p,q}\frac{\langle E_{p},E_{q}^{(0)}|(K_{L}-K_{R}^{*})^{2}|E_{m},E_{n}^{(0)}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle\\ &=-i\gamma\sum_{p}\frac{\langle E_{p}|K^{2}|E_{m}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pn}^{(0)}}\,|E_{p},E_{n}^{(0)}\rangle-i\gamma\sum_{q}\frac{\langle E_{q}|K^{2}|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{mq}^{(0)}}\,|E_{m},E_{q}^{(0)}\rangle\\ &\quad+2i\gamma\sum_{p,q}\frac{\langle E_{p}|K|E_{m}\rangle\,\langle E_{q}|K|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle.\end{split}$ (39)
Crucially, this functional dependence does not rely on $(m,n)$ being the lowest excited pair: the same perturbative structure holds for any unperturbed eigenmode. As a result, the constraints we derive on the trace norm for rapid mixing apply broadly across the spectrum, rather than being tied to the lowest excited state.
An analogous statement holds in the strong-dissipation limit: the trace-norm requirements obtained from the corresponding perturbation theory do not single out the lowest excited sector. Therefore, using the lowest excited state as a proxy in the rewritten mixing-time expression is sufficient for formulating and comparing rapid-mixing conditions in both weak- and strong-dissipation regimes.
Appendix C Appendix C: Constraints on the lowest excited state in the strong- and weak-dissipation regimes
In this appendix, we prove sufficient conditions under which the trace norm of the lowest excited state, $\mathrm{Tr}\,|\sigma_{1}|=\|\sigma_{1}\|_{1}$, is bounded by $\mathrm{poly}(L)$ in both the strong- and weak-dissipation regimes, as stated in Eqs. (51) and (68). This bound is the trace-norm requirement for rapid mixing, as formulated in Theorem 2 in the main text.
Throughout, the total Hilbert-space dimension is $N=d^{L}$. We use $\|\cdot\|_{1}$ for the trace norm and $\|\cdot\|_{2}$ for the Frobenius norm, $\|X\|_{2}^{2}=\mathrm{Tr}(X^{\dagger}X)$. For the lowest excited state operator $\sigma_{1}$ we assume the normalization $\mathrm{Tr}(\sigma_{1}^{2})=1$ (Hermitian $\sigma_{1}$), which implies $\|\sigma_{1}\|_{2}=1$.
C.1 Strong-dissipation regime
In the strong-dissipation regime, the unperturbed doubled-space (vectorized) eigenbasis is $\{|\epsilon_{p},\epsilon_{q}\rangle\}\equiv\{|\epsilon_{p}\rangle_{L}\otimes|\epsilon_{q}\rangle_{R}\}$, where $|\epsilon_{s}\rangle$ are simultaneous eigenstates of the (Hermitian) Lindblad operators $\{K_{j}\}$. The leading correction to the state $|\epsilon_{m},\epsilon_{m}\rangle$ induced by the Hamiltonian superoperator
$H_{s}\;=\;H_{L}\otimes\mathbbm{1}_{R}\;-\;\mathbbm{1}_{L}\otimes H_{R}^{T}$ (40)
takes the standard first-order perturbative form
$\begin{split}|\epsilon_{m},\epsilon_{m}^{(1)}\rangle&=\sum_{p,q}\frac{\langle\epsilon_{p},\epsilon_{q}|H_{s}|\epsilon_{m},\epsilon_{m}\rangle}{\epsilon^{(0)}_{mm}-\epsilon^{(0)}_{pq}}\,|\epsilon_{p},\epsilon_{q}\rangle\\ &=-\sum_{p}\frac{H_{pm}}{\epsilon^{(0)}_{pm}}\,|\epsilon_{p},\epsilon_{m}^{(0)}\rangle+\sum_{q}\frac{H_{qm}}{\epsilon^{(0)}_{mq}}\,|\epsilon_{m},\epsilon_{q}^{(0)}\rangle,\end{split}$ (41)
where $H_{pm}\equiv\langle\epsilon_{p}|H|\epsilon_{m}\rangle$, and $\epsilon^{(0)}_{pq}$ denotes the unperturbed doubled-space spectrum.
It is convenient to write the perturbed vectorized state as
$|\epsilon_{m},\epsilon_{m}^{(1)}\rangle=\sum_{p,q}C_{pq}\,|\epsilon_{p},\epsilon_{q}^{(0)}\rangle,$ (42)
which corresponds to an operator (matrix) representation
$\sigma_{1}=\sum_{p,q}C_{pq}\,|\epsilon_{p}\rangle\langle\epsilon_{q}|.$ (43)
From Eq. (41), the coefficients are
$C_{pq}=\frac{H_{qm}}{\epsilon^{(0)}_{mq}}\delta_{pm}-\frac{H_{pm}}{\epsilon^{(0)}_{pm}}\delta_{qm},\qquad(p\neq q).$ (44)
Equation (44) shows that $\sigma_{1}$ has at most $2N$ nonzero matrix elements in the $\{|E_{s}\rangle\}$ basis: only the $m$-th row and $m$-th column can be populated.
Cutoff decomposition.
Fix a cutoff $c>0$. Let $A$ be the matrix obtained from $\sigma_{1}$ by keeping only those entries with magnitude larger than $c$, and let $B:=\sigma_{1}-A$ be the remainder. Thus $\sigma_{1}=A+B$ and by the triangle inequality
$\|\sigma_{1}\|_{1}\leq\|A\|_{1}+\|B\|_{1}.$ (45)
We bound $\|A\|_{1}$ and $\|B\|_{1}$ separately.
Bound on the “large” part $A$.
Using $\|X\|_{1}\leq\sqrt{\mathrm{rank}(X)}\,\|X\|_{2}$ and $\|A\|_{2}\leq\|\sigma_{1}\|_{2}=1$,
$\|A\|_{1}\leq\sqrt{\mathrm{rank}(A)}\,\|A\|_{2}\leq\sqrt{\mathrm{rank}(A)}\leq\sqrt{N_{\mathrm{large}}},$ (46)
where $N_{\mathrm{large}}$ is the number of entries of $\sigma_{1}$ with magnitude >c>c, and we used $\mathrm{rank}(A)\leq N_{\mathrm{large}}$. Therefore, a sufficient condition for $\|A\|_{1}\leq\mathrm{poly}(L)$ is
$N_{\mathrm{large}}\leq\mathrm{poly}(L).$ (47)
Bound on the “small” part $B$.
Again using $\|X\|_{1}\leq\sqrt{\mathrm{rank}(X)}\,\|X\|_{2}$ and $\mathrm{rank}(B)\leq N$, we get
$\|B\|_{1}\leq\sqrt{N}\,\|B\|_{2}.$ (48)
Moreover, since $B$ contains only entries of magnitude $\leq c$ and $\sigma_{1}$ has at most $2N$ nonzero entries in this basis, we have the Frobenius bound
$\|B\|_{2}^{2}=\sum_{i,j}|B_{ij}|^{2}\leq(2N)\,c^{2},\qquad\Rightarrow\qquad\|B\|_{1}\leq\sqrt{N}\,\sqrt{2N}\,c\;=\;\sqrt{2}\,N\,c.$ (49)
Hence a sufficient condition for $\|B\|_{1}\leq\mathrm{poly}(L)$ is
$c\leq\frac{\mathrm{poly}(L)}{N}.$ (50)
From cutoff bounds to a sparsity condition on $H$.
From Eq. (44), the only potentially nonzero entries of $\sigma_{1}$ are proportional to $|H_{pm}|/|\epsilon_{pm}^{(0)}|$ and $|H_{qm}|/|\epsilon_{mq}^{(0)}|$. In the strong-dissipation regime, the denominators are set by the dissipative scale and remain bounded below by an $L$-independent constant (up to model-dependent factors), so bounding entries of $\sigma_{1}$ reduces to bounding the matrix elements of $H$ in the $\{|\epsilon_{s}\rangle\}$ basis. Choosing a cutoff that scales as $c\sim e^{-\alpha L}$ leads to the following sufficient condition:
$\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|\epsilon_{s}\rangle,\qquad N(\alpha,H,|\epsilon_{s}\rangle)\leq\mathrm{poly}(L),$ (51)
where $N(\alpha,H,|\epsilon_{s}\rangle)$ counts the number of indices $p$ such that
$|H_{sp}|\geq e^{-\alpha L}\qquad\text{for a fixed eigenstate }|\epsilon_{s}\rangle\text{ of }\{K_{j}\}.$ (52)
Together with Eqs. (47) and (50), this implies $\|\sigma_{1}\|_{1}=\mathrm{Tr}|\sigma_{1}|\leq\mathrm{poly}(L)$.
Other excited states.
For aribitary excited states, its first-order correction can be written as
$\begin{split}|\epsilon_{m},\epsilon_{n}^{(1)}\rangle&=\sum_{p,q}\frac{\langle\epsilon_{p},\epsilon_{q}|H_{s}|\epsilon_{m},\epsilon_{n}\rangle}{\epsilon^{(0)}_{mn}-\epsilon^{(0)}_{pq}}\,|\epsilon_{p},\epsilon_{q}\rangle\\ &=\sum_{p}\frac{H_{pm}}{\epsilon^{(0)}_{mn}-\epsilon^{(0)}_{pn}}\,|\epsilon_{p},\epsilon_{n}^{(0)}\rangle+\sum_{q}\frac{H_{qn}}{\epsilon^{(0)}_{mn}-\epsilon^{(0)}_{mq}}\,|\epsilon_{m},\epsilon_{q}^{(0)}\rangle.\end{split}$ (53)
Following a similar analysis, we find that the sufficient condition for the trace norm of an arbitrary excited state to be bounded by $\mathrm{poly}(L)$ is identical to that of the lowest excited state. Thus, the sparsity constrains for rapid mixing applies universally across the entire spectrum, rather than being restricted to the lowest-lying states.
C.2 Weak-dissipation regime
In the weak-dissipation regime, the leading correction to doubled-space eigenstates is generated by the dissipator. Let $\{|E_{p},E_{q}\rangle\}$ be the doubled-space basis built from energy eigenstates $\{|E_{s}\rangle\}$ of the Hamiltonian, and let $|E_{m},E_{n}^{(0)}\rangle$ denote an unperturbed eigenstate with eigenvalue $\epsilon_{mn}^{(0)}$. To first order in the dissipation strength $\gamma$, standard (non-degenerate) perturbation theory gives
$\begin{split}|E_{m},E_{n}^{(1)}\rangle&=-i\gamma\sum_{p,q}\frac{\langle E_{p},E_{q}^{(0)}|(K_{L}-K_{R}^{*})^{2}|E_{m},E_{n}^{(0)}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle\\ &=-i\gamma\sum_{p}\frac{\langle E_{p}|K^{2}|E_{m}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pn}^{(0)}}\,|E_{p},E_{n}^{(0)}\rangle-i\gamma\sum_{q}\frac{\langle E_{q}|K^{2}|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{mq}^{(0)}}\,|E_{m},E_{q}^{(0)}\rangle\\ &\quad+2i\gamma\sum_{p,q}\frac{\langle E_{p}|K|E_{m}\rangle\,\langle E_{q}|K|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle.\end{split}$ (54)
Here $K_{L}$ and $K_{R}$ denote left- and right-multiplication by $K$ in the doubled space, and we use the shorthand $K_{pm}\equiv\langle E_{p}|K|E_{m}\rangle$.
As in the strong-dissipation analysis, we unvectorize the perturbed doubled-space vector and view it as an operator $\sigma_{1}$, whose matrix elements are precisely the coefficients in Eq. (54). We bound $\|\sigma_{1}\|_{1}$ by splitting
$\sigma_{1}=A+B$ (55)
with a cutoff $c>0$: entries with magnitude $\geq c$ are assigned to $A$ (the “large” part), and the remaining entries to $B$ (the “small” part). By the triangle inequality (45), it suffices to control $\|A\|_{1}$ and $\|B\|_{1}$ separately.
Large part $A$.
The same counting argument as in the strong-dissipation case yields
$\|A\|_{1}\leq\sqrt{N_{\mathrm{large}}},$ (56)
where $N_{\mathrm{large}}$ is the number of matrix elements of $\sigma_{1}$ whose magnitude is at least $c$. Hence $\|A\|_{1}\leq\mathrm{poly}(L)$ follows from $N_{\mathrm{large}}\leq\mathrm{poly}(L)$.
Small part $B$.
Equation (54) contains two structurally distinct sets of coefficients:
  1. 1.
    terms proportional to $K_{pm}K_{qn}$ (the double sum in the last line), which populate $O(N^{2})$ entries in the $\{|E_{p},E_{q}\rangle\}$ basis;
  2. 2.
    terms proportional to $(K^{2})_{pm}$ or $(K^{2})_{qn}$ (the first two sums), which populate $O(N)$ entries.
Accordingly, we decompose
$B=B_{1}+B_{2},$ (57)
where $B_{1}$ collects the $O(N^{2})$ coefficients proportional to $K_{pm}K_{qn}$ and $B_{2}$ collects the $O(N)$ coefficients proportional to $(K^{2})_{pm}$ or $(K^{2})_{qn}$.
By construction, every entry of $B$ has magnitude at most $c$. Therefore,
$\|B_{1}\|_{2}^{2}\leq N^{2}c^{2}\qquad\Rightarrow\qquad\|B_{1}\|_{1}\leq\sqrt{N}\,\|B_{1}\|_{2}\leq N^{3/2}c,$ (58)
so $\|B_{1}\|_{1}\leq\mathrm{poly}(L)$ is ensured by
$c\leq\frac{\mathrm{poly}(L)}{N^{3/2}}.$ (59)
Similarly,
$\|B_{2}\|_{2}^{2}\leq Nc^{2}\qquad\Rightarrow\qquad\|B_{2}\|_{1}\leq\sqrt{N}\,\|B_{2}\|_{2}\leq Nc,$ (60)
and $\|B_{2}\|_{1}\leq\mathrm{poly}(L)$ follows from
$c\leq\frac{\mathrm{poly}(L)}{N}.$ (61)
From cutoff bounds to conditions on matrix elements of $K$.
To translate the cutoff constraints into a condition on $K$, we relate the typical size of the coefficients in $B_{1}$ and $B_{2}$ to the matrix elements $K_{pm}$.
For the $B_{1}$ sector, the relevant coefficients scale as products $K_{pm}K_{qn}$; hence a conservative estimate is
$|K_{pm}K_{qn}|\lesssim c\qquad\Rightarrow\qquad\max_{p,m}|K_{pm}|\lesssim\sqrt{c}.$ (62)
Combining with Eq. (59) yields the sufficient requirement
$\max_{p,m}|K_{pm}|\;\lesssim\;\frac{\mathrm{poly}(L)}{N^{3/4}}.$ (63)
For the $B_{2}$ sector, coefficients involve $(K^{2})_{pm}=\sum_{r}K_{pr}K_{rm}$. Using the bound
$|(K^{2})_{pm}|\leq\sum_{r}|K_{pr}||K_{rm}|\leq N\big(\max_{a,b}|K_{ab}|\big)^{2},$ (64)
we obtain
$|(K^{2})_{pm}|\lesssim c\qquad\Rightarrow\qquad\max_{p,m}|K_{pm}|\lesssim\sqrt{\frac{c}{N}}.$ (65)
Combining with Eq. (61) gives the sufficient requirement
$\max_{p,m}|K_{pm}|\;\lesssim\;\frac{\mathrm{poly}(L)}{N}.$ (66)
Since this is stronger than the $B_{1}$ condition for large $N$, it is sufficient to impose
$\max_{p,m}|K_{pm}|\;\lesssim\;\frac{\mathrm{poly}(L)}{N}.$ (67)
Equivalently, choosing an exponential cutoff $c\sim e^{-\alpha L}$ turns the above bounds into a sparsity requirement on $K$ in the energy eigenbasis. A sufficient condition is that there exists $\alpha\geq 1$ such that, for every eigenstate $|E_{s}\rangle$, only polynomially many matrix elements exceed the cutoff:
$\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|E_{s}\rangle,\qquad N(\alpha,K,|E_{s}\rangle)\leq\mathrm{poly}(L),$ (68)
where $N(\alpha,K,|E_{s}\rangle)$ counts the number of indices $p$ such that
$|K_{sp}|\geq e^{-\alpha L}.$ (69)
Combining the bounds on $A$ and $B$ with Eq. (68) and the cutoff choice $c\sim e^{-\alpha L}$ yields
$\|\sigma_{1}\|_{1}=\mathrm{Tr}|\sigma_{1}|\leq\mathrm{poly}(L)$ (70)
in the weak-dissipation regime.

### v2

@@
 Also, an estimation of the mixing time typically plays a crucial role in many of the quantum algorithmic applications to yield efficient quantum algorithms [29, 2].
-Also, mixing time is related to entropy [31, 46, 13], entanglement dynamics, and the structure of the Lindblad spectrum and the dissipative phase transition, where at the transition point, the mixing time diverges [46, 79, 45, 42, 57, 33, 32].
-Lots of progress has been made in analyzing the mixing time for the open system dynamics.
-Previous studies have shown that for Lindbladians satisfying quantum detailed balance, the system’s dynamics can be mapped to a local, frustration-free parent Hamiltonian in a doubled Hilbert space.
-The spectral gap of this Hamiltonian then provides a useful lower bound for the Lindbladian’s spectral gap, offering a powerful tool for estimating mixing time [15, 68, 35].
-A different approach uses the modified logarithmic Sobolev inequality [73, 9], which analyzes the exponential decay of relative entropy.
-This method can provide a tighter, polylog($N$) mixing time bound, offering a significant improvement over bounds based solely on the spectral gap [36, 11, 4, 3, 23].
-Besides, other research has estimated mixing time by analyzing the Hamiltonian and dissipative parts separately, although this is applicable in limited cases [20].
-Despite the recent progress, obtaining the mixing time of the Lindbladian dynamics, and analyzing its dependent of system size is typically difficult and can sometimes be the bottleneck of the analysis.
+Also, mixing time is related to entropy [32, 47, 14], entanglement dynamics, and the structure of the Lindblad spectrum [59, 60] and the dissipative phase transition, where at the transition point, the mixing time diverges [47, 82, 46, 43, 58, 34, 33].
+Considerable progress has been made in bounding mixing times for open quantum dynamics.
+For Lindbladians satisfying quantum detailed balance, the dynamics can be mapped to a local frustration-free parent Hamiltonian on a doubled Hilbert space, whose spectral gap lower-bounds the Lindbladian gap and enables mixing-time estimates [16, 71, 36].
+Complementarily, the modified logarithmic Sobolev inequality [76, 10] controls the exponential decay of relative entropy and can yield tighter, $\mathrm{polylog}(N)$ mixing-time bounds than gap-based estimates [37, 12, 5, 4, 24].
+Other works bound mixing by treating the Hamiltonian and dissipative parts separately, but only in restricted settings [21].
+Despite these advances, determining the mixing time and its system-size scaling for generic Lindbladians remains challenging and often becomes a bottleneck.
 In this letter, we map the definition of mixing time of the open quantum system to the distance between wavefunctions in a doubled Hilbert space, directly connecting its definition to the Liouvillian gap and the trace norm of the lowest excited state.
@@
 That is, $\tau_{\mathrm{mix}}(\eta)$ is the minimal time such that, for arbitrary initial state $\rho_{0}$, the evolved state $e^{\mathcal{L}t}[\rho_{0}]$ is within trace distance $\eta$ of the steady state $\sigma$.
-Since density matrices are Hermitian, this distance can also be written as
-$D(\rho(t),\sigma)=\frac{1}{2}\,\mathrm{Tr}\,|\rho(t)-\sigma|.$ (4)
+Since density matrices are Hermitian, this distance can be written as
+$D(\rho(t),\sigma)=\frac{1}{2}\,\mathrm{Tr}\,|\rho(t)-\sigma|.$
 We now introduce a mapping that relates the mixing time directly to spectral properties of the Lindbladian.
@@
 $\begin{split}&\rho=\sum_{m,n}\rho_{mn}\,|m\rangle\langle n|\\ \ \longmapsto\ &|\psi_{\rho}^{D}\rangle=\sum_{m,n}\rho_{mn}\,|m\rangle_{L}\otimes|n\rangle_{R}.\end{split}$ (4)
-Here $L$ and $R$ denote the left and right copies, respectively. Under this mapping, the Lindblad master equation becomes a Schrödinger-like equation in doubled space [76],
+Here $L$ and $R$ denote the left and right copies. Under this mapping, the Lindblad master equation becomes a Schrödinger-like equation in doubled space [79],
 $i\,\partial_{t}|\psi_{\rho}^{D}(t)\rangle=H^{D}\,|\psi_{\rho}^{D}(t)\rangle,$ (5)
 with a generally non-Hermitian generator $H^{D}=H_{s}-iH_{d}$ given by
-$\begin{split}H_{s}=&H_{L}\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes H_{R}^{T},\\ H_{d}=&\gamma\sum_{\alpha}\Big[-2\,K_{\alpha,L}\otimes K_{\alpha,R}^{*}\\ &+(K_{\alpha}^{\dagger}K_{\alpha})_{L}\otimes\mathcal{I}_{R}+\mathcal{I}_{L}\otimes(K_{\alpha}^{\dagger}K_{\alpha})_{R}^{*}\Big].\end{split}$ (7)
+$\begin{split}H_{s}=&H_{L}\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes H_{R}^{T},\\ H_{d}=&\gamma\sum_{\alpha}\Big[-2\,K_{\alpha,L}\otimes K_{\alpha,R}^{*}\\ &+\left(K_{\alpha}^{\dagger}K_{\alpha}\right)_{L}\otimes\mathcal{I}_{R}+\mathcal{I}_{L}\otimes\left(K_{\alpha}^{\dagger}K_{\alpha}\right)_{R}^{*}\Big].\end{split}$ (6)
 Here $H_{L/R}$ acts as $H$ on the corresponding copy, $T$ denotes transpose, denotes complex conjugation, and $\mathcal{I}$ denotes the identity.
@@
 where $\epsilon_{j}=\alpha_{j}-i\beta_{j}$ are the eigenvalues of $H^{D}$ (the Lindblad spectrum).
-The real parts $\alpha_{j}$ generate oscillations, while $\beta_{j}\geq 0$ set decay rates.
-[^1] Complete positivity implies $\beta_{j}\geq 0$.
+The real parts $\alpha_{j}$ generate oscillations, while $\beta_{j}\geq 0$ set decay rates  [^1] Complete positivity implies $\beta_{j}\geq 0$.
 Trace preservation fixes the steady mode: $c_{0}=1$, $\mathrm{Tr}\,\sigma_{0}=1$, and $\mathrm{Tr}\,\sigma_{j}=0$ for $j\neq 0$.
-Hermiticity preservation implies $\sigma_{j}=\sigma_{j}^{\dagger}$.
+Hermiticity preservation implies $\sigma_{j}=\sigma_{j}^{\dagger}$..
 Consequently, the deviation from the steady state $\sigma_{0}$ is
-$|\psi_{\rho}^{D}(t)\rangle-|\psi_{\sigma_{0}}\rangle=\sum_{j\neq 0}c_{j}\,e^{-i\alpha_{j}t-\beta_{j}t}\,|\psi_{\sigma_{j}}\rangle,$ (10)
+$|\psi_{\rho}^{D}(t)\rangle-|\psi_{\sigma_{0}}\rangle=\sum_{j\neq 0}c_{j}\,e^{-i\alpha_{j}t-\beta_{j}t}\,|\psi_{\sigma_{j}}\rangle,$
 where $|\psi_{\sigma_{0}}\rangle$ is the steady-state mode and $\{|\psi_{\sigma_{j}}\rangle\}_{j>0}$ are the decaying modes.
@@
 A convenient initial state with nonzero overlap on the slowest mode is
-$|\psi_{\rho_{0}}\rangle=|\psi_{\sigma_{0}}\rangle+c_{1}|{\psi_{\sigma_{1}}}\rangle,$ (11)
-where $\sigma_{0}$ is the steady state and $\sigma_{1}$ is the slowest decaying mode of the Lindbladian with eigenvalue $\Delta+i\alpha_{1}$ ($\Delta>0$).
-One might wonder whether the coefficient $c_{1}$ could grow with system size.
-This is ruled out by a purity bound: under vectorization, $\|\lvert\psi_{\rho_{0}}\rangle\|^{2}=\mathrm{Tr}(\rho_{0}^{2})\leq 1$, so the overlap (and hence $c_{1}$) is at most $O(1)$.
-We may therefore treat $c_{1}$ as an $O(1)$ constant.
-At time $t$,
-$|\psi_{\rho_{0}}(t)\rangle-|\psi_{\sigma_{0}}\rangle=c_{1}e^{-(\Delta-i\alpha_{1})t}\,|\psi_{\sigma_{1}}\rangle.$ (12)
-Consequently, the trace distance to the steady state is
-$D(\rho(t),\sigma_{0})=\frac{|c_{1}|}{2}\operatorname{Tr}|e^{-i\alpha_{1}t-\Delta t}\sigma_{1}|=\frac{|c_{1}|}{2}e^{-\Delta t}\operatorname{Tr}|\sigma_{1}|.$ (13)
+[…]
+[…]
+At time $t$, $|\psi_{\rho_{0}}(t)\rangle-|\psi_{\sigma_{0}}\rangle=c_{1}e^{-(\Delta-i\alpha_{1})t}\,|\psi_{\sigma_{1}}\rangle$. Consequently, the trace distance to the steady state is
+$D(\rho(t),\sigma_{0})=\frac{|c_{1}|}{2}e^{-\Delta t}\operatorname{Tr}|\sigma_{1}|.$ (10)
 The mixing time is thus
-$\tau_{\textit{mix}}(\eta)=\Delta^{-1}\left[\log(\operatorname{Tr}|\sigma_{1}|)-\log{(\frac{2\eta}{|c_{1}|})}\right].$ (14)
+$\tau_{\textit{mix}}(\eta)=\Delta^{-1}\left[\log\left(\operatorname{Tr}|\sigma_{1}|\right)-\log{\left(\frac{2\eta}{|c_{1}|}\right)}\right].$ (11)
 This reformulation of the mixing time indicates that the mixing time of an open system is set by the inverse Liouvillian gap, multiplied by a logarithmic factor controlled by the trace norm of the lowest excited state of the doubled-space Hamiltonian.
 One might worry that higher excited states, despite their larger decay rates, could have sufficiently smaller trace norms to yield a larger overall mixing-time bound.
-We explain in Appendix C why it is nevertheless sufficient to focus on the lowest excited state  [^2] We focus on the lowest excited state for two reasons.
+We explain in SM [1] why it is nevertheless sufficient to focus on the lowest excited state  [^2] We focus on the lowest excited state for two reasons.
 First, the fast mixing condition imposes no explicit constraints on trace norms across the spectrum, so the lowest excited state provides a natural and minimal diagnostic.
 Second, in both the weak- and strong-dissipation regimes, the rapid mixing requirement on the trace norm takes the same functional form for all excited states, and is therefore not specific to the lowest excited state.
-See Appendix C for details..
+See SM [1] for details..
 The trace norm of the lowest excited state has some explicit constraints. Using the Cauchy-Schwarz inequality, we have:
@@
 Combining this with the definition of mixing time, we arrive at a general upper bound of mixing time:
-$\tau_{\textit{mix}}(\eta)\leq\frac{1}{2}\Delta^{-1}\left[\log(N)-2\log{(\frac{2\eta}{|c_{1}|})}\right].$ (17)
+$\tau_{\textit{mix}}(\eta)\leq\frac{1}{2}\Delta^{-1}\left[\log(N)-2\log{\left(\frac{2\eta}{|c_{1}|}\right)}\right].$ (14)
 For a system of size $L$ and local Hilbet space dimension $d_{0}$,
-$\tau_{\text{mix}}(\eta)\leq\tfrac{1}{2}\Delta^{-1}\left[L\log d_{0}-2\log{(\frac{2\eta}{|c_{1}|})}\right].$ (18)
-This bound of mixing time is general but not always saturated.
-For example, non-Hermitian skin effects can localize eigenstates near boundaries [29], reducing effective support and giving $\tau_{\text{mix}}\sim\Delta^{-1}\text{poly}(\log L)$ for total system size $L$.
-Mixing time is therefore governed jointly by the Liouvillian gap and the spatial structure of eigenstates.
+$\tau_{\text{mix}}(\eta)\leq\tfrac{1}{2}\Delta^{-1}\left[L\log d_{0}-2\log{\left(\frac{2\eta}{|c_{1}|}\right)}\right].$ (15)
+This mixing-time bound is general but need not be tight: its prefactor depends on the trace norm of the slowest mode, which can vary widely with the spatial structure of Liouvillian eigenmodes.
+If the slow mode is effectively supported on a subextensive region, $\operatorname{Tr}|\sigma_{1}|$ may scale subexponentially (or even $\mathrm{poly}(L)$), reducing the estimate from the generic scaling implied by $\operatorname{Tr}|\sigma_{1}|\sim e^{cL}$ to much smaller forms.
+Non-Hermitian skin effects that accumulate eigenstates near boundaries [30] offer a plausible route to such reduced support; we leave an explicit Lindbladian construction and a quantitative verification of the trace-norm scaling to future work.
+Overall, mixing is governed jointly by the Liouvillian gap and the spatial structure of the slow eigenmodes.
 Requirements for fast and rapid mixing.— We first recall two standard notions of efficient convergence to the steady state.
@@
 Table 1: Conditions for fast and rapid mixing in terms of the inverse Liouvillian gap $\Delta^{-1}$ and the trace norm of the lowest excited state $\mathrm{Tr}|\sigma_{1}|$.
-We next specialize these general criteria to the weak- and strong-dissipation regimes, which yield concrete model-dependent sufficient conditions.
+We next apply these general criteria to the weak- and strong-dissipation regimes, where they simplify to explicit, checkable sufficient conditions expressed in terms of the Hamiltonian and Lindblad operators.
 The resulting rapid mixing conditions are summarized in Table 2.
@@
 For a fixed eigenstate $|E_{s}\rangle$ of $H$, $N(\alpha,K,|E_{s}\rangle)$ counts the number of indices $p$ with $|K_{sp}|\geq e^{-\alpha L}$.
-Strong-dissipation regime— In this regime, the dissipative part of the Liouvillian dominates, and the Hamiltonian (Schrödinger) contribution can be treated perturbatively.
+Strong-dissipation regime— In this regime, the Hamiltonian (Schrödinger) contribution can be treated perturbatively.
 We use degenerate perturbation theory to estimate the Liouvillian gap and the mixing time, and obtain the following fast-mixing corollary:
@@
 We characterize the corresponding matrix elements by typical scales,
-$\langle\psi_{0k}|V^{D}|\psi_{1\alpha}\rangle\sim J,\qquad\langle\psi_{0k}|V^{D}|\psi_{0l}\rangle\sim J^{\prime},$
-where $\{|\psi_{0k}\rangle\}$ spans the degenerate zeroth-order manifold (chosen to diagonalize $V^{D}$ within that subspace) and $\{|\psi_{1\alpha}\rangle\}$ denotes zeroth-order excited modes.
+$\langle\psi_{0k}^{L}|V^{D}|\psi_{1\alpha}^{R}\rangle\sim J,\qquad\langle\psi_{0k}^{L}|V^{D}|\psi_{0l}^{R}\rangle\sim J^{\prime},$
+where $\{|\psi_{0k}^{L}\rangle\}$ spans the degenerate zeroth-order manifold (chosen to diagonalize $V^{D}$ within that subspace) and $\{|\psi_{1\alpha}\rangle\}$ denotes zeroth-order excited modes.
 The leading nonzero decay rate of the perturbed steady manifold arises at second order via virtual transitions into the excited sector.
@@
 $E_{0k}\approx E^{(0)}_{0k}+V^{D}_{0k,0k}+\sum_{q}\frac{|V^{D}_{0k,1q}|^{2}}{E^{(0)}_{0k}-E^{(0)}_{1q}}$. (19)
-The diagonal shift $V^{D}_{0k,0k}$ contributes only to the real part, while the Liouvillian gap is set by the dissipative contribution from the second-order term,
-$\Delta\equiv\mathrm{Im}\,E_{0k}\sim\sum_{q}\frac{|V^{D}_{0k,1q}|^{2}}{E^{(0)}_{1q}-E^{(0)}_{0k}}\sim c\,\frac{J^{2}}{\gamma_{\mathrm{loc}}},$ (23)
-with an $\mathcal{O}(1)$ constant $c$ (for our boundary coupling structure, c=8c=8).
-[^4] The factor $8$ counts: (i) action on the left or right copy, (ii) action at the left or right boundary, and (iii) raising or lowering the boundary state, giving $2^{3}=8$ possibilities.
+Since the smallest diagonal shift $V^{D}_{0k,0k}$ is zero  [^4] In the vectorized representation, one has $V^{D}=H\otimes\mathbb{I}-\mathbb{I}\otimes H^{T}$, which acts as the commutator superoperator, $V^{D}\,|\rho\rangle\rangle=|[H,\rho]\rangle\rangle$.
+Therefore, any operator $\rho$ that commutes with $H$ is a zero mode of $V^{D}$.
+In particular, the identity operator (and more generally any function of $H$, or any $|m\rangle\langle m|$ in the energy basis) satisfies $H,\rho]=0$, implying an exact eigenvalue $\lambda=0$ of $V^{D}$.
+Since our $\{|\psi_{0k}^{R}\rangle,\langle\psi_{0k}^{L}|\}$ are chosen to diagonalize $V^{D}$ within the zeroth-order degenerate manifold, the diagonal matrix elements $\langle\psi_{0k}^{L}|V^{D}|\psi_{0k}^{R}\rangle$ coincide with these eigenvalues; hence the smallest value by magnitude is necessarily $\min_{k}|\lambda_{k}|=0$., thus the Liouvillian gap is set by the dissipative contribution from the second-order term,
+$\Delta\sim\sum_{q}\frac{|V^{D}_{0k,1q}|^{2}}{E^{(0)}_{1q}-E^{(0)}_{0k}}\sim c\,\frac{J^{2}}{\gamma_{\mathrm{loc}}},$ (20)
+with an $\mathcal{O}(1)$ constant $c$ (for our boundary coupling structure, c=8c=8) [^5] The factor $8$ counts: (i) action on the left or right copy, (ii) action at the left or right boundary, and (iii) raising or lowering the boundary state, giving $2^{3}=8$ possibilities..
 Crucially, $\Delta$ does not shrink with $L$: only Hamiltonian terms involving boundary sites connect the steady manifold to decaying modes, so the number of contributing excited states is set by the boundary size, i.e., $\mathcal{O}(1)$ in 1D.
 Using $\tau_{\mathrm{mix}}(\eta)\sim\Delta^{-1}\log\!\bigl(\mathrm{Tr}|\sigma_{1}|\bigr)$ (up to the standard $\eta$-dependent additive constant), we obtain
-$\tau_{\mathrm{mix}}(\eta)\sim\frac{\gamma_{\mathrm{loc}}}{c\,J^{2}}\log\!\bigl(\mathrm{Tr}|\sigma_{1}|\bigr).$ (24)
+$\tau_{\mathrm{mix}}(\eta)\sim\frac{\gamma_{\mathrm{loc}}}{c\,J^{2}}\log\!\bigl(\mathrm{Tr}|\sigma_{1}|\bigr).$
 Moreover, the first-order correction to the lowest excited mode generically spreads over the entire degenerate zeroth-order manifold. Since that manifold contains $d_{0}^{2L-2}$ basis states,
 $\log\!\bigl(\mathrm{Tr}|\sigma_{1}|\bigr)\leq\log\!\bigl(\sqrt{d_{0}^{2L-2}}\bigr)=(L-1)\log d_{0},$
-and this bound is saturated here because the $V^{D}$-diagonal basis typically forms superpositions across the full manifold. Therefore,
-$\tau_{\mathrm{mix}}(\eta)\sim\frac{\gamma_{\mathrm{loc}}}{c\,J^{2}}(\log d_{0})\,L,$ (25)
-which is polynomial in $L$ and hence implies fast mixing. In the Appendix we present the analogous analysis for strong bulk dissipation and show that this perturbative mechanism does not apply there.
+and this bound is saturated here because the $V^{D}$-diagonal basis typically forms superpositions across the full manifold.
+Therefore, $\tau_{\mathrm{mix}}(\eta)\sim\frac{\gamma_{\mathrm{loc}}}{c\,J^{2}}(\log d_{0})\,L$, which is $\mathrm{poly}(L)$ and hence implies fast mixing.
+In the SM [1], we analyze the strong bulk-dissipation regime and show that this perturbative mechanism does not apply.
 Physically, boundary-driven fast mixing is a quantum-Zeno-type stabilization: strong edge dissipation sets an $\mathcal{O}(1)$ scale, while boundary Hamiltonian couplings generate a size-independent gap $\Delta\propto J^{2}/\gamma_{\mathrm{loc}}$.
-Thus, boundary dissipation can guarantee scalable mixing even for complex bulk dynamics, whereas bulk dissipation typically requires different control assumptions (see Appendix).
-Finally, in the strong-dissipation regime, rapid mixing also requires control of the trace norm $\textrm{Tr}|\sigma_{1}|$, as stated in Theorem 2.
-To make this requirement explicit, we compute the leading correction to lowest excited doubled-space eigenstate in the vectorized $K$-eigenbasis $\{|\epsilon_{p},\epsilon_{q}\rangle\}$,
+Thus, boundary dissipation can guarantee scalable mixing even for complex bulk dynamics, whereas bulk dissipation typically requires different control assumptions (see SM [1]).
+Finally, in the strong-dissipation regime, rapid mixing also requires control of the trace norm $\mathrm{Tr}\,|\sigma_{1}|$, as stated in Theorem 2.
+To make this requirement explicit, we assume that $K$ is diagonalizable and compute the leading correction to the lowest excited doubled-space eigenstate in the vectorized $K$-eigenbasis $\{|\epsilon_{p},\epsilon_{q}\rangle\}$,
 $|\epsilon_{m},\epsilon_{m}^{(1)}\rangle=-\sum_{p}\frac{H_{pm}}{\epsilon^{(0)}_{pm}}\,|\epsilon_{p},\epsilon_{m}^{(0)}\rangle+\sum_{q}\frac{H_{qm}}{\epsilon^{(0)}_{mq}}\,|\epsilon_{m},\epsilon_{q}^{(0)}\rangle,$ (21)
-where $|\epsilon_{m}\rangle$ is the eigenstate of Lindblad operator $K$, $H_{pm}\equiv\langle\epsilon_{p}|H|\epsilon_{m}\rangle$, and $\epsilon^{(0)}_{pm}$ are the corresponding unperturbed doubled-space energy denominators.
+where $\{|\epsilon_{m}\rangle\}$ is the eigenstate of Lindblad operator $K$, $H_{pm}\equiv\langle\epsilon_{p}|H|\epsilon_{m}\rangle$, and $\epsilon^{(0)}_{pm}$ are the corresponding unperturbed doubled-space energy denominators.
 Requiring $\textrm{Tr}|\sigma_{1}|\leq\mathrm{poly}(L)$ then translates into a sparsity condition on $H$ in the $\{K_{j}\}$ eigenbasis:
-$\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|\epsilon_{s}\rangle,\qquad N(\alpha,H,|\epsilon_{s}\rangle)\leq\mathrm{poly}(L).$ (27)
-Here $N(\alpha,H,|\epsilon_{s}\rangle)$ counts the number of indices $p$ for which $|H_{sp}|\geq e^{-\alpha L}$, with $|\epsilon_{s}\rangle$ a fixed eigenstate of $\{K_{j}\}$.
-The proof is given in the Appendix.
-Weak-dissipation regime.— In the weak-dissipation regime we treat the dissipator as a perturbation to the unitary (Schrödinger) dynamics.
+$\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|\epsilon_{s}\rangle,\qquad N(\alpha,H,|\epsilon_{s}\rangle)\leq\mathrm{poly}(L).$
+Here $N(\alpha,H,|\epsilon_{s}\rangle)$ counts the number of indices $p$ for which $|H_{sp}|\geq e^{-\alpha L}$ (with $s$ fixed). The proof is given in the SM [1].
+Weak-dissipation regime—In the weak-dissipation regime, we treat the dissipator as a perturbation to the unitary (Schrödinger) dynamics.
 We estimate (i) the Liouvillian gap from the leading dissipative correction to the unperturbed spectrum and (ii) the trace norm of the lowest excited state from the corresponding first-order correction to the eigenvectors.
 Combining these ingredients yields the mixing-time scaling.
-Liouvillian gap. Consider an unperturbed doubled-space eigenstate
+Liouvillian gap. Consider an unperturbed doubled-space eigenstate with eigenvalue $\epsilon_{mn}^{(0)}$:
 $|E_{m},E_{n}^{(0)}\rangle\equiv|E_{m}\rangle_{L}\otimes|E_{n}\rangle_{R},\qquad\epsilon_{mn}^{(0)}=E_{m}-E_{n},$
-where $\{|E_{m}\rangle\}$ are eigenstates of $H$. The leading (first-order) correction from the dissipative term is
-$\langle E_{m},E_{n}^{(0)}|H_{d}|E_{m},E_{n}^{(0)}\rangle=-i\gamma\Bigl[\delta K_{m}^{2}+\delta K_{n}^{2}+\bigl(\overline{K_{m}}-\overline{K_{n}}\bigr)^{2}\Bigr],$ (28)
-with $\overline{K_{m}}\equiv\langle E_{m}|K|E_{m}\rangle$ and
-$\delta K_{m}^{2}\equiv\langle E_{m}|K^{2}|E_{m}\rangle-\overline{K_{m}}^{\,2}.$
+where $\{|E_{m}\rangle\}$ are the eigenstates of $H$. The first-order correction from the dissipative term is
+$\langle E_{m},E_{n}^{(0)}|H_{d}|E_{m},E_{n}^{(0)}\rangle=-i\gamma\Bigl[\delta K_{m}^{2}+\delta K_{n}^{2}+\bigl(\overline{K_{m}}-\overline{K_{n}}\bigr)^{2}\Bigr]$
+with $\overline{K_{m}}\equiv\langle E_{m}|K|E_{m}\rangle$ and $\delta K_{m}^{2}\equiv\langle E_{m}|K^{2}|E_{m}\rangle-\overline{K_{m}}^{\,2}$.
 When $K,H]\neq 0$, typical fluctuations scale as $\delta K_{m}^{2}\sim L^{a}$ with $a\geq 0$.
-For gapped systems one expects $a=0$, giving a Liouvillian gap of order $\mathcal{O}(\gamma)$; for gapless systems ($a>0$), enhanced fluctuations can parametrically reduce the gap.
-[^5] If $[K,H]=0$, then $\delta K_{m}^{2}=0$ and decay rates are set by the smallest nonzero value of $(\overline{K_{m}}-\overline{K_{n}})^{2}$.
-For local $K$ in an ETH-satisfying system this scale is typically $\mathcal{O}(1)$, yielding a size-independent gap.
+For gapped systems one expects $a=0$, giving a Liouvillian gap of order $\mathcal{O}(\gamma)$; for gapless systems ($a>0$), enhanced fluctuations can parametrically reduce the gap  [^6] If $[K,H]=0$, then $\delta K_{m}^{2}=0$ and decay rates are set by the smallest nonzero value of $(\overline{K_{m}}-\overline{K_{n}})^{2}$.
+For local $K$ in an ETH-satisfying system, this scale is typically $\mathcal{O}(1)$, yielding a size-independent gap..
 Therefore, a sufficient condition for rapid mixing based on the Liouvillian gap is that $H$ is gapped, which ensures a constant gap.
-Trace norm of the lowest excited state. The first-order correction to the corresponding excited eigenvector takes the form
-$\begin{split}|E_{m},E_{n}^{(1)}\rangle=-i\gamma\sum_{p,q,s,t}\frac{\bigl(K_{ps}-K_{qt}\bigr)\bigl(K_{sm}-K_{tn}\bigr)}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle,\end{split}$ (29)
-with $K_{ps}\equiv\langle E_{p}|K|E_{s}\rangle$.
-To ensure that the trace norm of the lowest excited state grows at most polynomially with system size, it is sufficient to impose a sparsity constraint on the matrix elements of $K$ in the $\{|E_{s}\rangle\}$ basis:
-$\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|E_{s}\rangle,\qquad N(\alpha,K,|E_{s}\rangle)\leq\mathrm{poly}(L),$ (30)
-where $N(\alpha,K,|E_{s}\rangle)$ counts the number of indices $p$ such that
-$|K_{sp}|\geq e^{-\alpha L}\quad\text{for fixed }|E_{s}\rangle.$
-A simple limiting example is the nearest-neighbor quantum Ising model (with Hamiltonian containing only $\sigma_{j}^{z}\sigma_{j+1}^{z}$ terms), with Lindblad operators taken to be arbitrary single-site operators.
-In this case the Hamiltonian eigenbasis is a product basis, and each local $K$ connects only $\mathcal{O}(1)$ configurations.
-Consequently one may take $N(\alpha,K,|E_{s}\rangle)=\mathcal{O}(1)$ (e.g., $N=1$ in the strictly diagonal case), illustrating the required sparsity.
-Discussion— By using the double-space mapping, we directly connected the mixing time to the Liouvillian gap and the trace norm of the lowest excited states.
+Trace norm of the lowest excited state. The first-order correction to the corresponding excited eigenvector is
+$\begin{split}|E_{m},E_{n}^{(1)}\rangle=-i\gamma\sum_{pqst}\frac{\bigl(K_{ps}-K_{qt}\bigr)\bigl(K_{sm}-K_{tn}\bigr)}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle,\end{split}$ (22)
+[…]
+[…]
+[…]
+Discussion— In this letter, by using the double-space mapping, we directly connected the mixing time to the Liouvillian gap and the trace norm of the lowest excited states.
 This connection provides a clear requirement for achieving fast and rapid mixing.
@@
 Our analysis shows that this is because localization properties directly affect the trace norm of the eigenstates, which in turn influences the mixing time.
-We analyzed the mixing time in both the strong and weak dissipation regimes, where the Liouvillian gap and trace norm can be calculated straightforwardly using perturbation theory.
+We analyze the mixing time in both the strong and weak dissipation regimes, where the Liouvillian gap and trace norm can be calculated straightforwardly using perturbation theory.
 This analysis not only simplifies the problem but also reveals general conditions for fast and rapid mixing.
-These findings are crucial for assessing the efficiency of quantum algorithms and state preparation using dissipative evolution.
-The general conditions we identified can be used to design specific forms of the Lindblad operator to more efficiently prepare desired quantum states.
-This work opens several interesting avenues for future research.
-First, it is important to compute the Liouvillian gap and the trace norm of the lowest excited states beyond the weak- and strong-dissipation limits, where our perturbative approach breaks down.
-Second, in connection with rapid mixing, it would be useful to identify conditions under which the trace norm of the lowest excited states is at most polynomial in system size, and to clarify how such conditions relate to features such as the locality of dissipation or the correlation length.
-A further question is whether changes in the scaling of the Liouvillian gap are tied to changes in the scaling of the lowest-mode trace norm.
-By tuning model parameters, one can compare the crossover where the inverse Liouvillian gap changes its system-size dependence (e.g., from polynomial to logarithmic) with the crossover where the trace norm changes (e.g., from exponential to polynomial).
-Determining whether these crossovers coincide—or how far apart they occur—could provide a new way to distinguish relaxation mechanisms.
-The boundary-driven transverse-field Ising chain, with dissipation only at the ends, provides a concrete setting in which both scalings vary with dissipation strength and can be directly compared.
-Finally, for dissipative state preparation, rapid mixing is essential to reach the target state on experimentally relevant timescales; the sparsity constraints derived here offer a practical guide for designing protocols that remain experimentally feasible.
-Acknowledgements— We thank Jerome Lloyd, Tian-Gang Zhou and Pengfei Zhang for helpful discussions.
+This work suggests several directions for future research.
+First, motivated by rapid mixing, it would be valuable to identify general conditions under which the trace norm of the slowest mode is bounded by $\mathrm{poly}(L)$, and to understand how these conditions relate to physical features such as locality of dissipation and correlation length.
+A closely related question is whether changes in the scaling of the Liouvillian gap are accompanied by changes in the scaling of the slowest-mode trace norm.
+Concretely, one can tune model parameters to compare (i) the crossover where the inverse gap changes its system-size dependence (e.g., from polynomial to logarithmic) with (ii) the crossover where the trace norm changes (e.g., from exponential to polynomial).
+Whether these crossovers coincide could provide a diagnostic of distinct relaxation mechanisms.
+The boundary-driven transverse-field Ising chain, with dissipation only at the ends, offers a concrete setting where both scalings vary with dissipation strength and can be directly compared.
+Finally, rapid mixing is crucial for dissipative state preparation on experimentally relevant timescales.
+Our sparsity constraints provide practical guidance for designing experimentally feasible protocols and for assessing the efficiency of dissipative algorithms.
+More broadly, the general conditions identified here can be leveraged to engineer Lindblad operators that prepare desired target states more efficiently.
+Acknowledgements— We thank Jerome Lloyd, Tian-Gang Zhou, Pengfei Zhang, Yu-Min Hu, Shuo Liu, and Yuxuan Zhang for helpful discussions.
+We are also grateful to Lorenza Viola and Yikang Zhang for their careful reading of the manuscript and for providing valuable feedback and suggestions.
 Appendix A Appendix A: Mixing time analysis for strong bulk dissipation
@@
 The corresponding zeroth-order generator can then be written as
-$H^{0,D}=-i\gamma\sum_{\alpha=1}^{L}(K_{\alpha,L}\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes K_{\alpha,R}^{*})^{2}.$ (31)
+$H^{0,D}=-i\gamma\sum_{\alpha=1}^{L}\left(K_{\alpha,L}\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes K_{\alpha,R}^{*}\right)^{2}.$ (23)
 The ground state(GS) of the zeroth-order Hamiltonian is the state that satisfies
@@
 We denote the matrix elements of the perturbation as $\langle\psi_{0k}|V^{D}|\psi_{1q}\rangle=J$, where $|\psi_{0k}\rangle$ represents a GS with label $k$ and $|\psi_{1q}\rangle$ represents a first excited state with label $q$.
-Since the diagonal (Hermitian) shift $V^{D}_{0k,0k}$ only affects the real part, it does not affect the Liouvillian gap.
-The only contribution to the Liouvillian gap is given by the last term in the Eq.(36)
-$\Delta\propto-iL\frac{J^{2}}{\gamma}.$
+Since the smallest diagonal shift $V^{D}_{0k,0k}$ vanishes, the Liouvillian gap is not determined by the first-order (diagonal) correction from $V^{D}$, but instead by the dissipative contribution from the second-order term.
+The vanishing of the smallest diagonal shift follows directly from the structure of $V^{D}$. In the vectorized representation,
+$V^{D}=H\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes H^{T},$ (29)
+so that $V^{D}$ acts as the commutator superoperator, $V^{D}\,|\rho\rangle\rangle=|[H,\rho]\rangle\rangle$.
+Hence any operator $\rho$ that commutes with $H$ is a zero mode of $V^{D}$.
+In particular, the identity operator $\mathcal{I}$, any function $f(H)$, and the projectors $|\ket{m}\bra{m}$ in the energy eigenbasis satisfy $H,\rho]=0$, which implies that $V^{D}$ has an exact eigenvalue $\lambda=0$.
+[…]
+[…]
+coincide with the corresponding eigenvalues of $V^{D}$ restricted to this manifold, and consequently the smallest value (in magnitude) is
+$\min_{k}|\lambda_{k}|=0.$ (31)
+It follows that the Liouvillian gap is controlled solely by the second-order term in Eq. (28), yielding the scaling
+$\Delta\propto-\,i\,L\,\frac{J^{2}}{\gamma}.$ (32)
 Because this energy correction diverges as $L\to\infty$, this perturbation analysis is not applicable in the thermodynamic limit.
@@
 As we derived in the main text, we arrive at:
-$\tau_{\textit{mix}}(\eta)=\Delta^{-1}\left[\log(\operatorname{Tr}|\sigma_{1}|)-\log{(\frac{2\eta}{|c_{1}|})}\right].$ (37)
+$\tau_{\textit{mix}}(\eta)=\Delta^{-1}\left[\log\left(\operatorname{Tr}|\sigma_{1}|\right)-\log{\left(\frac{2\eta}{|c_{1}|}\right)}\right].$ (33)
 Here, $\Delta$ is the Liouvillian gap and $\sigma_{1}$ denotes an lowest excited eigenmode (written as an operator in the original Hilbert space, or equivalently as a vector in the doubled space).
 This reformulation shows that the mixing time of an open system can be expressed as the inverse Liouvillian gap, multiplied by a logarithmic factor set by the trace norm of an excited state in the doubled (vectorized) space.
-A potential concern is that higher excited states, while decaying faster, might have substantially smaller trace norms and could in principle yield a larger estimate of the mixing time.
+A potential concern is that higher excited states, while decaying faster, might have substantially smaller trace norms and could, in principle, yield a larger estimate of the mixing time.
 For example, one could imagine that for some mode $(\beta_{2},\sigma_{2})$,
-$\beta_{2}^{-1}\log(\mathrm{Tr}|\sigma_{2}|)\;>\;\Delta^{-1}\log(\mathrm{Tr}|\sigma_{1}|),$ (38)
+$\beta_{2}^{-1}\log\left(\mathrm{Tr}|\sigma_{2}|\right)\;>\;\Delta^{-1}\log\left(\mathrm{Tr}|\sigma_{1}|\right),$ (34)
 so that including $\sigma_{2}$ would appear to increase the bound. Nevertheless, throughout this work we use the lowest excited state as a representative, for the following reasons.
@@
 To illustrate this in the weak-dissipation limit, the first-order correction to an arbitrary unperturbed eigenstate $|E_{m},E_{n}^{(0)}\rangle$ takes the form
-$\begin{split}|E_{m},E_{n}^{(1)}\rangle&=-i\gamma\sum_{p,q}\frac{\langle E_{p},E_{q}^{(0)}|(K_{L}-K_{R}^{*})^{2}|E_{m},E_{n}^{(0)}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle\\ &=-i\gamma\sum_{p}\frac{\langle E_{p}|K^{2}|E_{m}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pn}^{(0)}}\,|E_{p},E_{n}^{(0)}\rangle-i\gamma\sum_{q}\frac{\langle E_{q}|K^{2}|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{mq}^{(0)}}\,|E_{m},E_{q}^{(0)}\rangle\\ &\quad+2i\gamma\sum_{p,q}\frac{\langle E_{p}|K|E_{m}\rangle\,\langle E_{q}|K|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle.\end{split}$ (39)
+$\begin{split}|E_{m},E_{n}^{(1)}\rangle&=-i\gamma\sum_{p,q}\frac{\langle E_{p},E_{q}^{(0)}|\left(K_{L}-K_{R}^{*}\right)^{2}|E_{m},E_{n}^{(0)}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle\\ &=-i\gamma\sum_{p}\frac{\langle E_{p}|K^{2}|E_{m}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pn}^{(0)}}\,|E_{p},E_{n}^{(0)}\rangle-i\gamma\sum_{q}\frac{\langle E_{q}|K^{2}|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{mq}^{(0)}}\,|E_{m},E_{q}^{(0)}\rangle\\ &\quad+2i\gamma\sum_{p,q}\frac{\langle E_{p}|K|E_{m}\rangle\,\langle E_{q}|K|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle.\end{split}$ (35)
 Crucially, this functional dependence does not rely on $(m,n)$ being the lowest excited pair: the same perturbative structure holds for any unperturbed eigenmode.
@@
 C.1 Strong-dissipation regime
-In the strong-dissipation regime, the unperturbed doubled-space (vectorized) eigenbasis is $\{|\epsilon_{p},\epsilon_{q}\rangle\}\equiv\{|\epsilon_{p}\rangle_{L}\otimes|\epsilon_{q}\rangle_{R}\}$, where $|\epsilon_{s}\rangle$ are simultaneous eigenstates of the (Hermitian) Lindblad operators $\{K_{j}\}$.
-The leading correction to the state $|\epsilon_{m},\epsilon_{m}\rangle$ induced by the Hamiltonian superoperator
+In the strong-dissipation regime, we assume that $K$ is diagonalizable.
+The unperturbed doubled-space (vectorized) eigenbasis is $\{|\epsilon_{p},\epsilon_{q}\rangle\}\equiv\{|\epsilon_{p}\rangle_{L}\otimes|\epsilon_{q}\rangle_{R}\}$, where $|\epsilon_{s}\rangle$ are simultaneous eigenstates of the Lindblad operators $\{K_{j}\}$.
+The leading correction to the state $|\epsilon_{m},\epsilon_{m}\rangle$, induced by the Hamiltonian superoperator
 $H_{s}\;=\;H_{L}\otimes\mathbbm{1}_{R}\;-\;\mathbbm{1}_{L}\otimes H_{R}^{T}$ (36)
@@
 $\begin{split}|\epsilon_{m},\epsilon_{m}^{(1)}\rangle&=\sum_{p,q}\frac{\langle\epsilon_{p},\epsilon_{q}|H_{s}|\epsilon_{m},\epsilon_{m}\rangle}{\epsilon^{(0)}_{mm}-\epsilon^{(0)}_{pq}}\,|\epsilon_{p},\epsilon_{q}\rangle\\ &=-\sum_{p}\frac{H_{pm}}{\epsilon^{(0)}_{pm}}\,|\epsilon_{p},\epsilon_{m}^{(0)}\rangle+\sum_{q}\frac{H_{qm}}{\epsilon^{(0)}_{mq}}\,|\epsilon_{m},\epsilon_{q}^{(0)}\rangle,\end{split}$ (37)
-where $H_{pm}\equiv\langle\epsilon_{p}|H|\epsilon_{m}\rangle$, and $\epsilon^{(0)}_{pq}$ denotes the unperturbed doubled-space spectrum.
+where $H_{pm}\equiv\langle\epsilon_{p}|H|\epsilon_{m}\rangle$, and $\epsilon^{(0)}_{pq}$ denotes the unperturbed doubled-space eigenvalue associated with the basis state $|\epsilon_{p},\epsilon_{q}\rangle$.
+Here, $\langle\epsilon_{p}|$ denotes the $p$th left eigenvector of $K$.
+For a general non-Hermitian $K$, the left eigenstate is not necessarily the Hermitian (complex-conjugate) transpose of the corresponding right eigenstate $|\epsilon_{p}\rangle$.
 It is convenient to write the perturbed vectorized state as
@@
 To first order in the dissipation strength $\gamma$, standard (non-degenerate) perturbation theory gives
-$\begin{split}|E_{m},E_{n}^{(1)}\rangle&=-i\gamma\sum_{p,q}\frac{\langle E_{p},E_{q}^{(0)}|(K_{L}-K_{R}^{*})^{2}|E_{m},E_{n}^{(0)}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle\\ &=-i\gamma\sum_{p}\frac{\langle E_{p}|K^{2}|E_{m}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pn}^{(0)}}\,|E_{p},E_{n}^{(0)}\rangle-i\gamma\sum_{q}\frac{\langle E_{q}|K^{2}|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{mq}^{(0)}}\,|E_{m},E_{q}^{(0)}\rangle\\ &\quad+2i\gamma\sum_{p,q}\frac{\langle E_{p}|K|E_{m}\rangle\,\langle E_{q}|K|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle.\end{split}$ (54)
+$\begin{split}|E_{m},E_{n}^{(1)}\rangle&=-i\gamma\sum_{p,q}\frac{\langle E_{p},E_{q}^{(0)}|\left(K_{L}-K_{R}^{*}\right)^{2}|E_{m},E_{n}^{(0)}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle\\ &=-i\gamma\sum_{p}\frac{\langle E_{p}|K^{2}|E_{m}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pn}^{(0)}}\,|E_{p},E_{n}^{(0)}\rangle-i\gamma\sum_{q}\frac{\langle E_{q}|K^{2}|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{mq}^{(0)}}\,|E_{m},E_{q}^{(0)}\rangle\\ &\quad+2i\gamma\sum_{p,q}\frac{\langle E_{p}|K|E_{m}\rangle\,\langle E_{q}|K|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle.\end{split}$ (50)
 Here $K_{L}$ and $K_{R}$ denote left- and right-multiplication by $K$ in the doubled space, and we use the shorthand $K_{pm}\equiv\langle E_{p}|K|E_{m}\rangle$.
@@
 in the weak-dissipation regime.
+Appendix D Appendix D: Two equivalent ways of rewriting the mixing time
+In this appendix, we clarify two equivalent conventions for rewriting the mixing time.
+The first is the convention adopted in the main text, while the second is commonly used in earlier works.
+The apparent discrepancy between the two is purely a matter of normalization (i.e., where one “stores” an extensive factor), and does not reflect any physical inconsistency.
+One example is the statement in Ref.
+[60] that the prefactor of the slowest decaying mode scales as $e^{L}$, whereas in our Eq.
+(11) of the main text the corresponding coefficient $c_{1}$ is at most $O(1)$.
+These statements are consistent because the product of the coefficient and the mode—which is what enters the physical trace distance—is invariant under rescaling of eigenmodes.
+D.1 Way 1 (main text): Hilbert–Schmidt normalization
+In the main text we normalize the slowest decaying (right) mode $\sigma_{1}$ by
+$\operatorname{Tr}(\sigma_{1}^{2})=1.$ (67)
+We then generally write the initial state as
+$\rho_{0}=\sigma_{0}+c_{1}\sigma_{1},$ (68)
+where $\sigma_{0}$ denotes the steady state.
+Because $\rho_{0}$ is a physical density matrix, its purity obeys $\operatorname{Tr}(\rho_{0}^{2})\leq 1$. Expanding and applying Cauchy–Schwarz,
+$\operatorname{Tr}(\rho_{0}^{2})$ $=\operatorname{Tr}(\sigma_{0}^{2})+2c_{1}\operatorname{Tr}(\sigma_{0}\sigma_{1})+c_{1}^{2}\operatorname{Tr}(\sigma_{1}^{2})$
+$\leq\operatorname{Tr}(\sigma_{0}^{2})+2|c_{1}|\sqrt{\operatorname{Tr}(\sigma_{0}^{2})\operatorname{Tr}(\sigma_{1}^{2})}+|c_{1}|^{2}\operatorname{Tr}(\sigma_{1}^{2})$
+$=\operatorname{Tr}(\sigma_{0}^{2})+2|c_{1}|\sqrt{\operatorname{Tr}(\sigma_{0}^{2})}+|c_{1}|^{2}$
+$\leq 1+2|c_{1}|+|c_{1}|^{2}=(1+|c_{1}|)^{2}.$ (69)
+In the last line we used $\operatorname{Tr}(\sigma_{0}^{2})\leq 1$ and $\sqrt{\operatorname{Tr}(\sigma_{0}^{2})}\leq 1$.
+Therefore, the physical constraint $\operatorname{Tr}(\rho_{0}^{2})\leq 1$ implies that $|c_{1}|$ cannot grow with system size; in particular, $c_{1}=O(1)$ under our normalization.
+With this convention, any potential $e^{L}$ enhancement in the contribution of the slowest mode must come from the trace norm of $\sigma_{1}$, not from $c_{1}$.
+D.2 Way 2 (common in the literature): trace-norm normalization
+Some references, for example [60, 30] instead normalize the slow mode so that its trace norm is $O(1)$, e.g.
+$\operatorname{Tr}|\tilde{\sigma}_{1}|=1,$ (70)
+which can be achieved by rescaling
+$\tilde{\sigma}_{1}=a\,\sigma_{1},\qquad\tilde{c}_{1}=c_{1}/a.$ (71)
+If $\operatorname{Tr}|\sigma_{1}|\sim e^{L}$ in our convention, then choosing $a\sim 1/\operatorname{Tr}|\sigma_{1}|$ makes $\operatorname{Tr}|\tilde{\sigma}_{1}|=O(1)$ and transfers the exponential factor into $\tilde{c}_{1}$.
+This is exactly how one can obtain an $e^{L}$-scaling coefficient in Ref.
+[60] while still being consistent with $c_{1}=O(1)$ in our convention.
+D.3 Invariant quantity and mixing time
+The trace distance to the steady state depends only on the product of the coefficient and the trace norm of the mode. Keeping only the slowest mode for clarity, the time-evolved contribution reads
+$\rho(t)-\sigma_{0}=c_{1}\,e^{-i\alpha_{1}t-\Delta t}\,\sigma_{1},$ (72)
+and hence
+$D(\rho(t),\sigma_{0})=\frac{1}{2}\operatorname{Tr}|\rho(t)-\sigma_{0}|=\frac{|c_{1}|}{2}\,e^{-\Delta t}\,\operatorname{Tr}|\sigma_{1}|.$ (73)
+Under the rescaling $\sigma_{1}\mapsto\tilde{\sigma}_{1}=a\sigma_{1}$ and $c_{1}\mapsto\tilde{c}_{1}=c_{1}/a$, the product $|c_{1}|\operatorname{Tr}|\sigma_{1}|$ is unchanged, and therefore the mixing time extracted from (73) is identical in both conventions.
+In other words, different normalizations merely redistribute the same size dependence between the coefficient and the eigen-operator, without changing any physical prediction.
+D.4 Remark on bi-orthogonality of left/right eigenmodes
+Finally, we emphasize that in Eq.
+(11) of the main text, $c_{1}$ is simply the coefficient in the operator decomposition of $\rho_{0}$.
+We do not identify it with a particular overlap with a left eigenmode, such as
+$\tilde{c}_{1}=\langle L_{1}|\rho_{0}\rangle,$ (74)
+nor do we impose a normalization condition like $\langle L_{1}|R_{1}\rangle=1$ for left/right eigenmodes of the Lindbladian.
+Such overlaps and bi-orthogonal normalizations can themselves carry strong system-size dependence for non-normal superoperators.
+Our bound on $c_{1}$ relies only on the basic physical constraint $\operatorname{Tr}(\rho_{0}^{2})\leq 1$ for a density matrix, together with the Hilbert–Schmidt normalization $\operatorname{Tr}(\sigma_{1}^{2})=1$ used in the main text.